Introduction: The Exposition of the Question of the Meaning of Being

The Necessity, Structure, and Priority of the Question of Being

0.1 The Necessity of an Explicit Retrieve of the Question of Being

The question of “being”:

1. It is not the highest category, but is rather prior to categorization.

2. It is not a circular question because “being” is not itself a being.

3. It is not self-evident because our understanding of “being” does nothing to answer the question it supposedly renders useless.

Indeed, the question of “being” must be reformulated and reunderstood to enable the construction of an answer.

0.2 The Formal Structure of the Question of Being

The question is seeking “being”, so it must be guided there by our factical but horizonless understanding of “being”. The question mustn’t pursue causes and origins of beings, since “being” is prior to beings, and not itself a being. A new conceptualization is needed.

How can the question find a horizonless concept? By understanding it in terms of the interrogator the question can see beings “interrogated with regard to their being”, see the being of the questioner.

The Da-sein, the “being with regard to its being”, is the question’s subject of interrogation. Is this circular? “A circle in reasoning cannot possibly lie in the formulation of the question…, because in answering this question it is not a matter of grounding by deduction but rather of laying bare and exhibiting the ground.” The Da-sein is where the factical understanding of “being” occurs.

0.3 The Ontological Priority of the Question of Being

Science’s crucial paradigm shifts are brought about by crises in their basic concepts. This refers to a reunderstanding of their fundamental principles, in a sense a reformation of their questions. Sciences arise from the domains of beings, so all of these questions rely on an investigation into the constitution of their being. Thus the investigation into “being” is prior to all scientific investigation. This priority must be understood as the fundamental task of ontology—clarifying the meaning of “being”.

0.4 The Ontic Priority of the Question of Being

Da-sein is ontically distinct in its ontological character. Understanding the question of existence ontically, through existence itself, is “existentiell” understanding. Understanding the structure of existence ontologically, through the “possibility and necessity in the ontic constitution of Da-sein”, is “existential”. Science, understanding Da-sein’s “being in a world”, requires an understanding of the being of the beings accessible in the world. Ontologies exploring these beings are rooted in Da-sein’s onticality. This brings us to the scene of our inquiry, where Da-sein always already understands “being”—its ontic structure. Fundamental ontology begins with an existential analysis of Da-sein.

Da-sein is prior to all other beings in the investigation due to its onticality, its ontological determination, and its ontic-ontological access to the ontologies of other beings in a world. Pre-ontological understanding of being is what Da-sein brings to the interpretation of being. It is not only the first being to look to in this interpretation, but also that being which is always already related to what is sought.

The Double Task in Working Out the Question of Being: The Method of the Investigation and Its Outline

0.5 The Ontological Analysis of Da-sein as the Exposure of the Horizon for an Interpretation of the Meaning of Being in General

How do we get proper access to the being of Da-sein? Though ontically it is immediately available, ontologically we must find Da-sein’s being through the world, since Da-sein is “essentially, continually, and most closely related to” the world. Da-sein must “show itself to itself on its own terms…initially and for the most part—in its average everydayness.” This project is not intended to give a complete analysis of Da-sein, but to know it enough to catch “being”, since that first piece of the puzzle is necessary for all ontological inquiry.

We must show how time is the horizon by which Da-sein understands being. This understanding of time and the common one will be shown to originate from “temporality”, “the being of Da-sein which understands being.” Time is commonly (naïvely) understood to distinguish different regions of beings. This understanding must be reevaluated.

The approach here, of understanding and conceiving being in terms of time, reveals being as having temporal character. Heidegger uses the term “temporal determination” to clarify this determination of the meaning of being.

0.6 The Task of a Destructuring of the History of Ontology

Da-sein grows according to its past—its future is its past in that it grows accustomed to an interpretation of itself and lives accordingly. Da-sein “occurs”, happens, in a manner enabling “world history”. The constitution of this happening is called historicity, Da-sein’s past, and it is made possible by temporality. Historical Da-sein, Da-sein’s inquiry into and pursuit of its past (“tradition”), leads us to temporality. Inquiry into being is characterized by historicity. The question must recognize and address its historicity and come into possession of its past. Da-sein is simultaneously entangled in the world and its own tradition, meaning Da-sein does not control the direction of its own inquiries and actions, particularly those concerning ontology. This tradition clouds the source of Da-sein’s existentiell understanding, and even that it exists or should be sought, even in the pursuit of fundamental principles. Greek ontology demonstrates how Da-sein, free of traditional distractions, understands being in terms of the world. As this became tradition the question was considered as having been answered by tradition, rendering Greek ontology as doctrine. Philosophy endures thanks to the soundness of its tradition’s genesis, but fails where it neglects the primacy of the genesis over the tradition. In terms of the question of being and its direction, we must take apart the tradition of ancient ontology. This is not to destroy tradition, but to give it the proper context. Kant’s investigation into time suffered from taking the meaning of being for granted and from its rooting in Descartes’ cogito. Without the world, the cogito’s “sum” has no way of establishing its meaning.

The Greeks moved from the dialectic, the replacement of Da-sein with “zōon logon echon”, to “legein” and “noein”, apprehension of an objectively present (in the world) being in its objective presence (in its “presencing”, its revealing of itself in its making present). “Ousia”—presence.

Okay, so destructure the tradition according to the question of being.

0.7 The Phenomenological Method of the Investigation

This ontology does not call itself “ontology” out of connection or obligation to a particular philosophical discipline; this ontology is prior to any other philosophical inquiry. The phenomenological treatment of the question does not indicate work within any doctrine, but a method of philosophical research. That phenomenology is so surely and clearly dedicated to working with “the things themselves” and not wacky metaphysics or arbitrary frameworks should put any such concerns to rest. We shall work out here a Greek meaning of “phenomenology” in terms of its components and its whole:

a: The Concept of Phenomenon—

• Phenomenon—“what shows itself in itself, what is manifest”

• Phenomena—“the totality of what lies in the light of day or can be brought to light”

• Phenomenon—“what looks like something, what ‘seems’, ‘semblance’ ”

• These definitions are connected, but we stick with the first one—it is not a semblance, but a being’s showing of itself in accordance with its meaning. This is not “appearance”! Symptoms are the appearance of an illness—they show themselves and the illness, but the illness does not show itself. Appearances are dependent upon phenomena.

b: The Concept of Logos—

• Logos—speech? But what does that mean?

• dēloun—making manifest

• apophainesthai—to let beings be unconcealed

c: The Preliminary Concept of Phenomenology—

• apophainesthai ta phainomena—“to let what shows itself be seen from itself, just as it shows itself from itself”

• Does this term, like “theology”, say what is being researched and what that contains? No. It simply says that its objects are held such that they are directly considered. Truth of being is concealed in that it does not show itself initially and for the most part, but it belongs to that which does “in such a way that it constitutes its meaning and ground.”

The phenomenological meaning of “phenomenon”, self-showing, means the being of beings. Phenomenology is necessary to encounter phenomena, and is thus a necessary part of ontology. It is necessary because of how phenomena can be occulted. One’s relation to a phenomenon can be one of ignorance, where it is unknown and not known to be unknown. A phenomenon can be known and then covered over, wholly or in part. When a previously discovered phenomenon is thus distorted one risks deception. This tends to occur whenever phenomena are formalized in a system, even one as ubiquitous as speech. To avoid this one must carefully secure an “ ‘originary’ and ‘intuitive’ grasp and explication of phenomena” instead of a mere observation.

Phenomenal—“what is given and is explicable in the way we encounter the phenomenon” Phenomenological—“everything that belongs to the manner of indication and explication, and constitutes the conceptual tools this research requires”

0.8 The Outline of the Treatise

We are trying to grasp being by first grasping Da-sein, finding there the horizon for understanding and interpreting being. First we will interpret Da-sein according to temporality, then show time as the transcendental horizon of the Quesion. Second we will begin destructuring the history of ontology.

The Interpretation of Da-sein in Terms of Temporality

Division One: The Preparatory Fundamental Analysis of Da-sein

The Exposition of the Task of a Preparatory Analysis of Da-sein

0.9 The Theme of the Analytic of Da-sein

By “existence” we do not mean “objective presence” but the determination of being of Da-sein. Since the “essence” of Da-sein lies in this meaning of “existence”, the way in which Da-sein is concerned with its being, it is nonsense to talk of its being in terms of objectively present attributes in its outward appearance but in terms of possible ways for it to be. “Da-sein” does not express a what, but a how. An objectively present being is indifferent to its own being. The possibility of Da-sein is never a feature or attribute of Da-sein, but is rather Da-sein itself. When Da-sein is “winning” itself in its being we call it “authentic”, but to be authentic or inauthentic (to win, lose, or ignore itself) is always the choice of Da-sein, a choice it has precisely because it is this possibility. The inauthentic moments are in no way “lesser”—they are some of the most important.

Its “existence” as here described and its always-being-mine (Heidegger stresses the importance of the first person) show how unique Da-sein is phenomenally and phenomenologically. Da-sein, in not being merely objectively present, presents itself in a manner so unusual that we must take the time to correctly present it.

Looking at the indifferent way in which Da-sein is initially and for the most part, we find a positive phenomenal characteristic of “averageness”. “Average everydayness” is Da-sein’s “ontic immediacy”. This ontically near realm is ontologically the farthest and most difficult to grasp. Existentiality is still here in average everydayness and inauthenticity, and they are thus not mere “aspects” but relations to its being.

Existentials —characteristics of being of Da-sein

Categories —determinations of being of beings unlike Da-sein. From “kategoresthai”—ontologically, to let a being be seen for everyone in its being. What are thus brought to light are the “kategoriai”.

“Existentials and categories are the two fundamental possibilities of the characteristics of being.” Their corresponding beings require different means of interrogation—who or what, respectively. Connecting the two is part of the task before us.

Now we must demonstrate the urgency of the existential analytic of Da-sein, its primacy over anthropology, psychology, and biology.

0.10 How the Analytic of Da-sein is to be Distinguished from Anthropology, Psychology, and Biology

Anthropology, psychology, and biology are today too distant from the ontological problematic. Modern thought was shaped by Descartes’ “cogito sum”, which expended much ink defining and determining the cogito without ever questioning the sum. More crucially, Descartes concerned himself with the cogitare of the “ego”, a subject fatally lacking the phenomenal content of Da-sein. Even when one considers a “subject” without asking the proper ontological questions, that subject carries the being of beings along underneath it, and no ontic questions will suffice until its thirst for ontology is satisfied. However we conceive of a being, we must keep ties to ontology. Scheler says a person is connected to acts—not objective presence, but “only to be experienced in the process itself and given in reflection.” Even this does not penetrate beyond the ontology of action into the being of Da-sein. Traditional (Christian) anthropology looks to some inadequate ontological foundations of personhood: the (still merely objectively present) rational animal and the construction of man in God’s image (saying that human being has roots in something greater). These Greek and theological definitions do not question being at all, and “the anthropological problematic remains undetermined in its decisive ontological foundation.”

Psychology looks to anthropology, and biology’s attempts to be a “science of life” root it in Da-sein’s ontology. “The ontology of life… determines what must be the case if there can be anything like just-being-alive.” Life is ontologically undetermined, and not a foundation for Da-sein.

Anthropology, psychology, and biology accomplish what they do without adequate ontology of Da-sein. We see that “hypotheses derived from empirical material” do not lead us to ontological foundations but distract from the ontological foundations always already “there” in the material. Claiming their self-evidence does nothing to address their importance.

0.11 The Existential Analytic and the Interpretation of Primitime Da-sein: The Difficulties in Securing a “Natural Concept of the World”

Primitive Da-sein can speak from a “more primordial absorption in ‘phenomena’.” Though it seems crude and simple it can be of great use. Primitive Da-sein is not everydayness, which is a kind of being of Dasein found in highly developed culture. Ethnology should not be looked to for an understanding of human being. Whatever positivistic disciplines or philosophy herself have done to develop a “natural concept of the world”, we have shown that this mere ordering is a distraction from the real problem of ontology, which will contribute indirectly to these disciplines but ultimately pursues its own goal.

Being-in-the-World in General as the Fundamental Constitution of Da-sein

0.12 A Preliminary Sketch of Being-in-the-World in Terms of the Orientation toward Being-in as Such

Mineness, the possibility of authenticity, inauthenticity, or “modal indifference” to these. These determinations of Da-sein’s being are grounded in being-in-the-world, a constitution of being which must be interpreted in order to know where to begin. While the term expresses a unified concept that can’t be broken down, we can look at it from different angles:

1. “In-the-world”—this perspective demonstrates the need for defining “worldliness” and determining the ontological structure of the world. (Chapter 3)

2. The “being”—that which (who, actually) is in the way of being-in-the-world. The average everydayness of Da-sein. (Chapter 4)

3. “Being in”—the ontological constitution of in-ness. (Chapter 5)

All three will have their time, but first we spend a little time on “being in”. We talk of being-in something, “as water is ‘in’ the glass”. But this spatial sort of in-ness applies to beings that are objectively present, related categorically. These beings are unlike Da-sein. For Da-sein, being-in expresses an existential. “Being” here is to dwell near and be familiar with …, and “Being-in is the formal existential expression of the being of Da-sein”, which is always being-in-the-world.

In terms of “being with”, to say that two beings touch, are somehow together or that one is encountered by the other, both can not be merely objectively present. A world in which beings can be encountered is necessary for a being to be accessible in its objective presence in the first place. The “worldless” objects cannot be together with each other. The way in which Da-sein has being-in-the-world and sees itself as together with the beings encountered in its world is called its facticity. Da-sein also has its own kind of objective presence, a “being-in-space” grounded always in its being-in-the-world. Da-sein has existential spaciality, in that its inness is not that of an object placed in space but of a factical being together with objects in the world.

Some ways of being-in refer to “taking care of” …, indicating being with a consideration of or apprehension about . . .. “Taking care” is an existential designating the being of a possible being-in-the-world. “Care” is to be understood ontologically, not as emotion. Da-sein is always “in the world”, it cannot be in any other way than “being-in”. Being-in is largely represented by knowing the world. Being-in-the-world’s connection with knowing the world must be examined as an existential of being-in.

0.13 The Exemplification of Being-in in a Founded Mode: Knowing the World

Knowledge is not a relation between subject and object, but a relation between Da-sein and world (and not in the same manner as that between subject and object!). We must thus characterize knowing as a phenomenon. That being which is known (nature) is not where the knowing is to be found, but in those beings which know. Knowing is neither objectively present nor apparent externally. “Knowing is a mode of being of Da-sein as being-in-the-world”, but then what is knowledge? “Being-in-the-world, as taking care of things, is taken in by the world which it takes care of.” Knowing determines what is objectively present, meaning there’s been a “deficiency” of taking care of the world and having to do with it. It is to look at a being encountered in its eidos (outward appearance), in a direction and from a perspective, and dwelling with it independently to have a perception of it. Looking at a being as something, one interprets and defines it, enabling its expression as a proposition. This is not the kind of proposition so separate from the world that we must ask as to its “correspondence”. This is not a “reaching out of” some inside for Da-sein, since Da-sein is always already outside. This knowing is Da-sein staying inside and out. Even knowledge forgotten or muddled is “a modification of primordial being-in”. Knowing is a new perspective and possibility of being, a mode of Da-sein. We must now look prior to being-in-the-world to understand it.

The Worldliness of the World

0.14 The Idea of the Worldliness of the World in General

We start an analysis and grounding for being-in-the-world by exploring the phenomenon of “world”—that which shows itself as being and the structure of being is a phenomenon, and we shall study the things “in” the world in these terms, “conceptually and categorically”. “Neither the ontic description of innerworldly beings nor the ontological interpretation of the being of these beings gets as such at the phenomenon of ‘world’. In both kinds of access to ‘objective being’, ‘world’ is already ‘presupposed’ in various ways.” Is “world” something subjective, something that is a character of being of Da-sein? But we are all in a world, the same world. This must be encuntered as the worldliness of world in general. “World” is a terms with many uses:

1. The totality of beings objectively present.

2. The being of beings or ranges of beings objectively present.

3. That “in which” a factical Da-sein lives.

4. Worldliness, that which can be modified into the respective structural totality of particular “worlds”.

The expression world refers to meaning 3. “World”, with quotes, refers to meaning 1. Traditional ontology has ignored the phenomenon of worldliness and being-in-the-world. It has instead tried to interpret the world in terms of nature, but ontologically beings can be discovered as nature only in a definite mode of being-in-the-world. Nature is so grounded in the concept of worldliness that it cannot hope to render it intelligible, since it is itself meaningless until worldliness is explained.

Average everydayness is the key. We will look at Da-sein’s nearest kind of being to find how Da-sein skips over worldliness to know the world.

The worldliness of the surrounding world (environmentality) will be sought through ontological analysis of the beings encountered therein.

The world “around” us takes on a spatial meaning, but this spatiality is not primary and arises out of the concept of worldliness. This analysis of worldliness will stand is educational contrast to that of Descartes.

A: Analysis of Environmentality and Worldliness in General—

0.15 The Being of Beings Encountered in the Surrounding World

Everyday being-in-the-world is association in the world with innerworldly beings. The closest association is “a handling, using, and taking care of things which has its own kind of ‘knowledge’.” By exploring phenomenologically the beings that show themselves in this taking care and looking to them with a phenomenological “knowing”, we hope to determine the structure of their being. This brings about the understanding of being always already possessed by Da-sein. We must look to “beings as they are encountered of their own accord in taking care,” and by rejecting “interpretational tendencies” we do not enter into “taking care” but go back to it. We look to “things” as our pre-phenomenal basis, as our beings.

But it is tempting to consider a “thing” in terms of thingliness or reality and ignore the pre-ontological way in which we take care of these things. We must look to useful things, and elucidate their kind of being. “Strictly speaking, there ‘is’ no such thing as a useful thing. There always belongs to the being of a useful thing a totality of useful things in which this useful thing can be what it is.” Useful means “in order to…”, which contains a reference of one thing to another. A room full of useful objects is not encountered as spatial relations but as “material for living”. The totality of useful things is encountered before any individual useful thing.

A “useful thing’s kind of being in which it reveals itself by itself” we shall call handiness. That useful things do not merely occur, but have this “being-in-themselves” that they are handy. To not just look at outward appearance, not just look at things theoretically, but to associate things in their “in-order-to” is circumspection.

Everyday association looks not to the useful things, the tools, but to the work, what is to be produced— the what-for. The work to be produced has the kind of being of a useful things. But following this backwards we find that not everything has been produced this way—in smoe useful things we discover “nature”, products of nature. A forest is timber, a mountain a quarry, a river is power, and wind is “the wind in the sails”. The work also points to the wearer/user of what is to be produced, “he ‘is’ there as the work emerges”. When a work is produced by the dozen, “beings with the kind of being of Da-sein for whom what is produced becomes handy” are encountered in the work, and along with them the world in which such beings live, our world. The work taken care of is at hand now in the domestic world, the public world, and the surrounding world of nature, since roads, bridges, rooftops and lights all associate with nature. This is not a mere spin, a subjective coloration of beings objectively present. If handiness turns out to be the kind of being of beings first discovered within the world, we do not get “world” simply by connecting the beings. How do we move from here to the phenomenon of world?

0.16 The Worldly Character of the Surrounding World Making ITself Known in Innerworldly Beings

World itself is prior to and necessary for innerworldly beings to be encountered and discovered. Does Da-sein’s encounter with useful things in the world point us to worldliness? A useful thing, in no longer associating itself properly with other useful things, becomes conspicuous in its unusability, unhandiness. The change it has undergone is not the change in qualities of a merely objectively present thing When a thing is missing, not at hand, unhandy, we suddenly find those things that are at hand obtrusive. Suddenly the objective presence of what is at hand is discovered when we find ourselves helpless to affect it, when it loses its association. So when objective presence makes itself known it does so alongside the handiness of useful things. The usefulness takes its leave in the form of conspicuousness, and handiness is shown alongside the worldly character of what is at hand. In a disruption of association we become ontically aware of a thing’s “what-for”, and the totality of usefulness makes itself known along with the world. When the world appears in the modes of taking care just interpreted, what is at hand loses its worldliness and appears as something merely objectively present.

Being-in-itself happens where there is inconspicuousness, unobtrusiveness, and nonobstinacy. But world cannot do this.

Being-in-the-world is circumspection, absorption in the references of handiness. Taking care of things works from a familiarity with the world. There are still questions, and we are still not ready to answer them.

0.17 Reference and Signs

Let us consider more deeply the phenomenon of reference. We wish to find a useful thing in terms of which references can be found, and that useful thing is a sign. Being-a-sign-for smoething can be formalized as a universal kind of relation. Looking to the sign itself as a useful thing, we find its usefulness consists in indicating or, very formally, relating.

Indicating is a subset of Reference is a subset of Relation.

Consider a turn signal—its usefulness is reference, indication. That is, its indicating can be taken as a kind of “referring”. This “referring” as indicating is not the ontological structure of the sign as a useful thing but rather the structure of being of useful things—serviceability for. Indicating’s “referral” is the “ontic concretion of the what-for of serviceability, and determines a useful thing for that what-for.” “Serviceability for” is an “ontological, categorical determination of the useful thing as useful thing.” The sign shows the distinction between referral as serviceability and referral an indicating. “Useful things which indicate have an eminent use in heedful association.” This must be clarified.

The meaning of a sign’s indication can only be ascertained by defining the appropriate way of associating with indicating things. We must comprehend their handiness. The being toward a turn signal is “yielding” or “standing still”, with reference toward the car. Da-sein is always “underway”, and so these ways of taking direction belong to Da-sein’s being-in-the-world, as they are boundary instances of its directed being. When we see the sign as an indicating thing that occurs, we do not “comprehend” it—comprehension is not circumspection. Circumspect overseeing “acquires an orientation with the surrounding world.” Signs give us access to the context of what is at hand such that heedful association secures an orientation. Signs don’t merely point from one thing to another but they bring a totality of useful things to circumspection along with the worldly character of what is at hand. Signs primarily indicate their own relevance.

Signs are established as useful things taking over the “work” of letting things at hand become conspicuous. Sometimes something already at hand is “taken as a sign”. “What is taken as a sign first becomes accessible through its handiness.” A farmer “accepting” the south wind as a sign of rain is not adding meaning to a objectively present being but is discovering the south wind in its being through circumspection, “by taking the lay of the land into account.” How is a sign grasped before it is established as a sign? Is it merely occurring or was one ignorant of its use up until now? Usefulness not discovered by circumspection is not the same as mere thingliness presented for the comprehension of something objectively present.

Signs’ conspicuousness “document the inconspicuousness constitutive for what is at hand nearest to us”. The sign “takes its conspicuousness from the inconspicuousness of the totality of useful things at hand in everydayness”. A string around one’s finger is narrowly intelligible and widely applicable, and when its established use is forgotten its maker, the only one for whom it is a sign, finds it obtrusive.

The coincidence of a sign with that which it indicates among primitive peoples means the sign has not become free from that for which it is a sign.

The relation between sign and reference:

1. The indicating is based upon the in-order-to (reference), as a possible concretion of the what-for of serviceability.

2. As the character of useful things at hand, the indicating of signs belongs to a referential context (the totality of useful things).

3. Signs make the surrounding world explicitly accessible to circumspection in their handiness.

Signs are ontically at hand, yet in their usefulness indicate the ontological structure of handiness, referential totality, and worldliness.

Reference is not comprehensible as a sign if it is ontologically the foundation for signs.

0.18 Relevance and Significance: The Worldliness of the World

Things at hand are at hand in terms of the world. How is the freeing of what is encountered within the world for circumspection the ontological distinction of the world?

Things at hand are suited or unsuited for things, and their “qualities” (like the “indicating” of a sign or the “hammering” of a hammer) are bound up with that suitability in the same way objective presence is with handiness.

That things at hand have the structure of reference means they have the character of being referred. Reference is reciprocal—beings are relevant together with other beings. Reference is “together…with…”.

Relevance is the being of innerworldly beings, and it is a relevance about the what-for of serviceability, which can itself be relevant. A hammer has to do with hammering has to do with fastening has to do with protection against weather] The protection “is” for the sake of a possibility of Da-sein’s being. Which relevance a thing has is prefigured in terms of the total relevance. The total relevance leads back to a what-for that has no relevance itself, it’s “not a being of the kind of being of things at hand within a world, but is a being whose being is defined as being-in-the-world, to whose constitution of being worldliness itself belongs.” This primary what-for isn’t just another link of relevance, but a “for-the-sake-of-which”. The for-the-sake-of-which always points to the being of Da-sein which is concerned about this being in its being. Before returning to this link from relevance to the being of Da-sein we must clarify “letting something be relevant”.

An “a priori” letting be relevant means to discover a previously existing being in its handiness and let it be encountered as the being of this being. This is a very ontic encounter with beings.

An ontological letting be relevant means freeing every thing at hand, ontically relevant or not, as a thing at hand and rather than “letting it be” as the discovered being it is work over it, improve it, destroy it. The previous freeing of beings for their innerworldly handiness. “The with-what…is freed in terms of the together-with-what of relevance.” When a being is discovered in its being, it is always already a thing at hand in the surrounding world rather than some kind of “stuff” objectively present and nothing more. Letting beings be relevant, freeing them for a totality of relevance, has already disclosed that for which it is freeing. That for which things at hand are freed, made accessible as innerworldly beings, is not discoverable if discoveredness refers only to a possibility of being of beings unlike Da-sein, since it (that for which this freeing happens) is exactly the understanding of world to which Da-sein is always already related.

Da-sein is referred by the context of relations to an in-order-to in terms of a potentiality-for-its-being, either authentic or inauthentic. Da-sein is always referred to the with-what of relevance, meaning it always already encounters beings as things at hand.

Da-sein lets beings be encountered for the context in which it understands itself.

The “wherein of self-referential understanding”, that for which one lets beings be encountered in the kind of being of relevance, is the phenomenon of world.

Worldliness is the structure of that to whch Da-sein is referred. Now we must understand the ontological context of Da-sein’s self-referral. Da-sein finds the context of things at hand along with its being—in its being Da-sein has alway already referred itself to an encounter with a “world”. This arises from a familiarity with significance, which gives Da-sein the possibility to understand and interpret, even to use language. I guess this makes sense is S.31. Understanding holds the relations here considered in “a preliminary disclosure”, and I’m going to start working this out on the next page.

Understanding—holds the relations of Da-sein to … in a preliminary disclosure. In its familiar beingin-relevance, understanding comes prior to the disclosure as that within which the reference moves. The relations themselves can concern understanding, and the relational character of such “referential relations” is signifying.

Da-sein “signifies” itself, gives itself to understand its being and potentiality-of-being. For-the-sake-of-which signifies in-order-to signifies what-for signifies what-in, letting something be relevant. These relations form a totality, the signifying in which Da-sein primordially understands its being-in-the-world. This “relational totality of signification” is called significance.

Da-sein is the ontic condition of the possibility of the discovery of beings in terms of their relevance in a world such that they can reveal themselves in their in-itself. This “familiarity with significance” means Da-sein, by being, is always referred to (together with) a world. Da-sein is essentially referred.

The significations with which Da-sein is familiar include the possible discovery of something life significations from itself, which lead to the possibility of words and language.

The concern now is that characterizing the being of what is at hand and worldliness as a referential context will jeapordize their “substantial being” by turning them into relations, which are always “something thought”. Does this turn the being of innerworldly beings into “pure thought”?

Here’s a look at the different structures and dimensions of the ontological problematic:

1. Handiness—the being of the innerworldly beings initially encountered.

2. Objective Presence—the being of beings that is found by discovering them in their own right in going through beings initially encountered.

3. The Worldliness of the World—the being of the ontic condition of the possibility of discovering innerworldly beings in general.

1 and 2 are categories concerning beings unlike Da-sein, but 3 is an existential determination of Da-sein, of being-in-the-world. The relations of the in-order-to, for-the-sake-of, with-what, etc. of relevance can’t be mathematically functionalized phenomenally, and they are not mere thoughts arising out of “thinking” but relations in which heedful circumspection dwells. Worldliness doesn’t volatize the being of innerworldly beings because they are only discoverable on the basis of worldliness. This discovery is the only way of making merely objectively present things accessible.

Now we’ll pause to piss on Descartes.

B: Contrast between Our Analysis of Worldliness and Descartes’ Interpretation of the World—Descartes’ interpretation of world begins with an innerworldly being and never looks at the phenomenon of world again. This extreme position considers the world as extensio, essentially the same as mere spatiality.

Section 19 will discuss the determination of the “world” as res extensa. Section 20 will discuss the foundations of this ontological determination. Section 21 is a hermeneutical look at Cartesian ontology of “world”.

0.19 The Determination of the “World” as Res Extensa

At the outset Descartes splits the ego cogito into res cogitans (“spirit”) and res corporea (“nature”). Descartes is the foundation of this messy ontology. He calls the being of beings in themselves substantia—substantiality is the being of beings as substance, particular substances are beings themselves.

he accesses substances through “attributes”, properties which essentially determine a substance’s substantiality. The res corporea is constituted by length, breadth, depth. “World” is spatial attributes. Extension must be there for any other determination of being to happen. Descartes figured that every other property of substance (division, shape, motion) is contingent on spatiality, extension. This was also explanation for how a corporeal being can change the distribution of its extension and present itself yet as the same being (since it is the dame underlying extension).

Force, color, hardness, etc. are excluded since if they were essential features then a being which shies away from the hand to evade touch and feeling would be denied its corporeal being. Extensio is what persists in a substance throughout all changes. Remaining constant. What remains constant characterizes the substantiality of the substance.

0.20 The Fundaments of the Ontological Definition of the “World”

Substance means a being in need of no other being in order to be. Whatever satisfies this idea in its truest sense, that which in its being needs no other being, is the ens perfectissimum, a purey ontological meaning for Descartes’ “Deus”. Everything else for him needs to be produced and sustained. Thus the production of what is objectively present and the lack of needing production are his horizon for understanding “being”. Everything other than God is ens creatum. Though both are “beings”, the gap between ens perfectissimum and ens creatum is supposedly “infinite”! How broad this understanding of being is! In light of this broadness, some created beings are called substances in that they need God for creation and sustenance but no other being of the “world”—they need no ens creatum. “There are two such substances: res cogitans and res extensa.” When we say “God is” and “the world is”, “is” cannot signify these beings in the same sense. If it were univocal, the same in both cases, their substantiality would be altered. The difference in meaning was considered “understood” and “self-evident”, and the question left unanswered. Descartes claimed that “No signification of this name (substance) which would be common to God and his creation can be distinctly understood.” He even goes so far as to say (and Kant goes so far as to repeat) that since being does not “affect” us it cannot be perceived. This dismisses the problematic of being entirely, and Descartes has to define substances some other way. So he says we must access beings not in terms of their being but their attributes, and thinking and extension are the attributes used to access res cogitans and res corporea, respectively.

Res extensa achieves priority thusly, through an evasion of substance’s “most substantial attribute”. Descartes muddies ontic and ontological meanings in the ambiguous term “substance” since he is busy with the meanings of words instead of working with “the things themselves”.

0.21 Hermeneutical Discussion of the Cartesian Ontology of the “World”

Descartes pretty much missed the point entirely. We looked for a kind of being before which provided access to the phenomenon of worldliness, so what kind of being of Da-sein has the being as extensio Descartes equates with that of the “world”? Mathematical knowing, as deductive, necessary knowledge, is in the true sense, it remains constant. The being of the “world” is handed to it by an idea, innerworldly beings do not present their own being but that of an ungrounded idea of being as constant objective presence. Mathematical knowledge is not his basis for ontology, but rather the kind of knowledge best suited to his ontological bias toward being as constant objective presence.

Descartes dismisses knowledge from senses as unimportant—it does not let us know any being in its being, but merely helps humans use and protect their bodies. True knowing comes through intuition, noein.

He determines hardness as an unyielding of one object before another! There is no consideration of the actual experience of hardness from that being, just a change or lack of change in spatial relationships.

Though Descartes values “intellective apprehension” above all other form of knowledge, by looking at being as constant objective presence he is unable to ontologically consider attitudes of this being, Da-sein, to other beings. Was Descartes even equipped to find a definite innerworldly being and its being with the world without having any understanding of the world, and thus innerworldliness? He wanted to raise the question, and he claimed to do so radically, yet he remained stuck in the mud of traditional ontology. In adding value based on his ontological standpoints he adds nothing, since in looking to their relation to the “fundamental stratum” of being for their value Descartes is only once again looking back to his ontological bias for the origin of “good things”. He can’t even find the concept of handiness through his idea of “valuable” qualities, since as close as he seems his love of independent, constant objective presence resists relation and circumspection! Now we see the structure of the critique of Cartesian ontology before us:

1. Where does the constant passing over of the phenomenon of world come from, and why is it so decisive for us?

2. Why do innerworldly beings take over for it as the ontological theme?

3. Why are they initially found in “nature”?

4. Why does the “rounding out” of this ontology happen through the phenomenon of value?

So Descartes may be fixable. Well, that was fun!

C: The Aroundness of the Surrounding World and the Spatiality of Da-sein—What role does spatiality play in Da-sein, being-in-the-world, innerworldly beings, and the world?

0.22 The Spatiality of Innerworldly Things at Hand

Innerworldly things at hand. We need to determine the spatiality of things, grasp it phenomenally and connect it to being.

When something is initially at hand, it is at hand first and it is also near. Handiness hints at this, but a thing can be near not by measuring distance but in terms of handling, use, and circumspection. Useful things are not just there in space, but are installed and set up. They are either in place or they lie around, and neither means they are in arbitrary spatial positions. Place isn’t a where but a definite there of a useful thing belonging there. This refers to a belonging in a totality of useful things. The whereto of a possible belonging somewhere is the region of a thing.

Region is not just direction toward but orbit around. The region must be discovered before any possibility is to be had of finding places. The orientation of places into regions constitutes the groundness of beings encountered initially in the surrounding world. The spatiality of what is at hand places these innerworldly beings for us, “where” a thing is is discovered in terms of everyday associations and not a catalogue of cartesian (coincidence?) coordinates.

The “surrounding world” doesn’t “settle in” to an existing space—its “worldliness articulates in its significance the relevant context of an actual totality of places circumspectly referred to each other.” The world discovers its own space!

That what is at hand can be encountered in its space of the surrounding world is possible only because Da-sein is “spatial” in its being-in-the-world.

0.23 The Spatiality of Being-in-the-World

Da-sein is not a being at hand in a place and is not in a position in “world space”. The spatiality of Da-sein’s being-in shows itself in de-distancing and directionality. De-distancing is making distance vanish. As de-distancing, Da-sein is letting beings be encountered in nearness. Remoteness and distance are only discovered through this de-distancing (an existential). Two things can not be distant from each other, since things are in their being incapable of de-distancing. The measurable distance between them is encountered by Da-sein in its de-distancing. Da-sein essentially brings things near, as our inventions indicate.

Estimated distances, though not explicitly precise, have their own definite meanings in terms of Da-sein’s everydayness. “Half an hour”, “a stone’s throw”, these statements make no attempt to measure but rather demonstrate their meaning in terms of heedful circumspection. Even if we hear an explicitly calculated “official” measurement we initially estimate it circumspectly (“it’s about three football fields long”). What is at hand isn’t there for Berkeley’s Big Brother, but for the circumspect, heedful everydayness of Dasein! Knowledge of objective distances is blind to the remoteness and nearness of what is at hand in the world. These distances are not “subjective” in terms of arbitrary conceptions of beings existing otherwise in themselves, but the discovery of beings with which Da-sein as existing is always already together in the world.

Da-sein, as always essentially de-distancing, associates in a slightly remote distance from itself. Spectacles, telephone receivers, and the road beneath one’s feet are all so near in their usefulness as to be almost totally inconspicuous (until they become obstacles). Da-sein is only “here” in that it exists primarily “over thre”, and comes back to itself.

Da-sein is spatial in that it encounters space by way of circumspection. It finds the distance “between” two things by de-distancing mhe distance. Da-sein cannot “cross over” de-distancing.

Every bringing near takes a direction in a region from what is de-distanced, so that its place may be found. This is the character of directionality in Da-sein. “Circumspect heedfulness is a directional de-distancing.” We now understand “signs” as things that take over and direct us. Directing and de-distancing are modes of being-in-the-world guided by the circumspection of heedfulness.

The feeling of right and left is not a “difference between two sides” but a searching for things. Left and right are directions, an orientation in a world. This directednes is grounded in directionality, which is grounded in Da-sein’s being-in-the-world. Kant’s worldless subject prevented him from understanding orientation in this way.

0.24 The Spatiality of Da-sein and Space

Being-in-the-world, in the sense of letting innerworldly beings be encountered, is “giving space”, making room, freeing things at hand for their spatiality. This is an existential of being-in-the-world. Through the spatiality thus encountered Da-sein discovers space. Da-sein is primordially spatial, and so space is encountered prior to encounters with other things in the surrounding world.

Space cannot be conceived as having the kind of being of res extensa, so it can’t be ontologically determined as a “phenomenon” of this res, and it is not merely “subjective” like the res cogitans.

Space isn’t the phenomenon of world, and it doesn’t show the primary ontological character of the being of innerworldly beings. To understand space we must go back to the world, find how space and the spatiality of Da-sein’s being-in-the-world constitute the world.

Being-in-the-World as Being-with and Being a Self: The “They”

Everyday Da-sein is taken in by its world. The phenomenon anwering to the question of “who?” is a mode of being of Da-sein.

0.25 The Approach to the Existential Question of the Who of Da-sein

Who is Da-sein? The who is that which stays the same as its mode of behavior and experiences change. That which maintains itself as identical. Da-sein has something which presents itself as objective presence. but this isn’t enough; this is the mode of being of a being unlike Da-sein.

Is the who of everyday Da-sein I myself? Is it simple perceiving reflection? We can say ontically that it is I with confidence, but ontologically we must exercise greater care and doubt. Also, we have already shown that an isolated I without a world does not exist initially, is never given. We must make visible the phenomenon of Mitda-sein, what is nearest in everydayness, and interpret it ontologically. We must not be misled by ontic characteristics when doing this.

So we shall interpret the “I” existentially! Da-sein is itself in existing, and existence (not objective presence) is the “substance” of human being.

0.26 The Mitda-sein of the Others and Everyday Being-with

We talked before about useful things pointing to other people. We encounter these others as those for whom the useful things are at hand. The world of Da-sein frees beings that are not handy or objectively present, but have the kind of being of Da-sein. They are like the Da-sein that frees them, there too and there with it.

These others are not distinguished from the I—they have the same kind of being, are most importantly those from whom one does not distinguish oneself; one is one! Being-there-too… Da-sein is with these others existentially. Not in terms of objective presence with but heedful circumspection with! The world is always this one that is shared, a with-world. Being-in is being-with, so the innerworldly being-in-itself of others is Mitda-sein.

It is phenomenal fact that others are encountered in the surrounding world. Da-sein finds itself in its doing, in what it needs, expects, and takes care of in the surrounding world. Da-sein understands itself in terms of its world, and this frequently includes Mitda-sein. Mitda-sein are not encountered as objectively present but “at work”, in their being-in-the-world. Even just standing around without heedfulness or circumspection is an existential, and is encountered as such.

Da-sein is essentially being-with. That is to say, being-with existentially determines Da-sein. It does this even without an other factically present. Being-alone is even just another way of being-with a deficient mode whose possibility proves being-with. Da-sein can be alone even when among many others, in that their Mitda-sein is encountered in the mode of an outsider.

Being-with is in Da-sein, encountered by another as Mitda-sein. Mitda-sein does not have the kind of being of a useful thing, but is itself Da-sein. It is not taken care of, but is instead a matter of concern. Initially and for the most part we live in deficient modes of concern, such as passing-one-another-by or not-mattering-to-one-another. This is the average everyday being-with-one-another, and it shows inconspicuousness and obviousness.

A positive concern can either take care of the other to the point of taking his “care” away from him (leaving him dependent, separate from the process, “inheriting” a finished product) or leap ahead of the other and return care to him. This latter takes care of the other authentically as Mitda-sein and the former as a what.

Concern is guided by considerateness and tolerance.

Da-sein is being-with, “is” for the sake of others. Others are disclosed together with the being-with, are a part of worldliness as the for-the-sake-of-which. This is how Mitda-sein shows itself in heedful being in terms of the things at hand in the world. Since the being of Da-sein is being-with, and so when Da-sein circumspectly finds and takes care of things it does so together with the other. The other is initially disclosed in the taking care of concern.

0.27 Everyday Being One’s Self and the They

“In what is taken care of in the surrounding world, the others are encountered as what they are; they are what they do.” Da-sein always compares itself to others, concerns itself with the distance between itself and them. Being-with-one-another has the character of distantiality. This existential places the everyday possibilities of being of Da-sein in others’ care. The others’ influence, the they (“das Man”), becomes more and more inconspicuous the more Da-sein is drawn into it. As we live like the they and relate more to it, we find the they more and more determining our everydayness. In this way, Da-sein is disburdened. Life is made easy by recourse to the they, which everyone is in everydayness but which is nobody. The they is an existential. The self of everyday Da-sein is the they-self, as opposed to the authentic self found by “breaking up the disguises with which Da-sein cuts itself off from itself.”

Authentic being one’s self is not separate from the they, but is an existentiell modification of its existentiality.

The everyday kind of being, absorbed in the world, passes over the phenomenon of world and replaces it with things, in their objective presence in the world. This is how the ontological interpretation of being has been passed over for so long. Ironically, the they (which is in no way objectively present) is the source of this misdirection. Authentic being thus does not happen by running or detaching from the they, since the authentic self has even less the character of objective presence, but is an existentiel modification of the existentiality of the they.

Being-in as Such

0.28 The Task of a Thematic Analysis of Being-in

It’s time to look back to being-in, both for a better understanding of b-i-t-w and to begin understanding Da-sein’s primordial being, care. We have already explained being-in-the-world as being together with the world (taking care of things), being-with (concern), and being one’s self (who). Rather than dissolving these into some single primordial relation, since they are all already primordial we shall recognize them as equiprimordial.

Being-in is not an expression of objective presence “in” another objectively present thing, as an objectively present “subject” in an objectively present “world”, but as a kind of being of Da-sein. The phenomenon of being-in is not a commercium between objectively present things, it is not even the “between” itself, since there is nothing to be between. We are not talking of an excluded middle, but a whole that we must take care not to split.

Da-sein is its there. “There” happens where a being has disclosed itself in its spatiality. Da-sein discloses itself and is “there” for itself together with the world. This disclosure is aletheia, revealing itself as beingin-the-world, to itself be the clearing. This is the only way for objectively present things to even become accessible. Da-sein is this disclosure, so we must explore and analyze being-in as disclosure as well as how Da-sein is its “there”, is its disclosure, in an everyday way.

There are two equiprimordially constitutive existentials of the there: attunement and understanding. These are part A of the chapter, and discourse, which determines them, is part B.

A: The Existential Constitution of the There—

0.29 Da-sein as Attunement

Attunement is the ontological side of what we understand ontically as being in a mood. Da-sein is always in a mood—that is to say, not being in a mood is when Da-sein is tired of itself, when being manifests itself as a burden. “Mood makes manifest ‘how one is and is coming along’ ”, bringing being to its “there”.

Da-sein is disclosed with its mood as being delivered to its being as what it has to be. This disclosure doesn’t mean to be known; Da-sein ontically and existentielly evades the being disclosed in moods. This evasion unveils Da-sein as being delivered over to the there—the there is disclosed in the evasion. This expresses the thrownness of Da-sein into its there. Da-sein is thrown into its there as being-in-the-world. Thrownness suggests the facticity of Da-sein’s being delivered over.

Attunement brings Da-sein before itself. Da-sein finds itself in attunement not by seekingbut by fleeing mood discloses in the mode of tuning away from thrownness. First essential ontological characteristic of attunemnt: “Attunement discloses Da-sein in its thrownness, initially and for the most part in the mode of an evasive thing turning away.” The there is disclosed in Da-sein’s evasion of the being disclosed in its mood.

This disclosure makes “experiences” possible. Mood discloses being-in-the-world and makes directedness possible. Second essential ontological characteristic of attunement: “It is a fundamental existential mode of being of the equiprimordial disclosedness of world, being-there-with, and existence because this disclosure itself is essentially being-in-the-world.”

Being-in is existentially determined such that what it encounters matters to it according to its attunement, which discloses the world as something one feels about in such and such a way. Only something attuned as fearing or fearless can encounter things at hand in terms of their threat.

This “mattering” that mood grants things is necessary even for sense perceptions. Attunement has a “disclosive submission” to world that allows things that matter to be encountered. This should not “surrender science ontically to ‘feeling’ ”, but demonstrate the role of directedness in scientific inquiry.

The phenomenological interpretation of Da-sein must let it interpret itself. We will demonstrate the phenomenon of attunement mor concretely as we find it in the definite mode of fear. We now interpret the most existential-ontologically significant attunement of Da-sein, Angst.

0.30 Fear as a Mode of Attunement

Three aspects of fear: what we fear, fearing, and why we fear. These belong together, and bring attunement into the picture. The fearsome thing is always a thing in the world, something with the kind of being of handiness, objective presence, or Mitda-sein. The fearsome thing has the character of being threatening. It (1)shows itself in a context of relevance as harmful, which is directed at a range of what can be affected. This means (2)it comes from a definite region, which (3)is and begets things that are “uncanny”. (4)Threat indicates that the harmful thing is not here yet, but is comng nearer. In coming nearer, (5)the sense that it may or may not cause harm intensifies. (6)The possibility that the harm never occurs does not reduce the fear, but intensifies it. Circumspection sees it as fearsome because it is in the attunement of fear. That is, fearfulness attumes the world as one in which things can harm us. Fear is always afraid of the fearful being itself, Da-sein. When we fear the loss of things, it is because “Da-sein is in terms of what it takes care of.” This what applies to Mitda-sein as we fear for someone. As we fear something happning “at any moment” we feel alarm. When what we fear is unfamiliar, we feel horror. Alarm and horrow together are terror. There are so many others, because Da-sein, if attuned in the mode of fearfulness, can encounter any situation with fear. Different attunements bring us to encounter the same situations differently.

0.31 Da-sein as Understanding

“World is ‘there’; its Da-sein is being-in. Being-in is ‘there’ as that for the sake of which Da-sein is.” The disclosure of being-in-the-world in this for-the-sake-of-which is understanding. “Significance is that for which world as such is disclosed.”

Ontically, understanding often means to grok, to be able to, to be ready to . . .. Used existentially, the … “is not a what, but being as existing.” Da-sein’s existing as a potentiality of being comes from this existential understanding. Da-sein’s ability to … is not tacked on to what it is, but rather it is being-possible. These possibilities are concerning taking care of the “world”, concern for others, and being itself, for its own sake. Attunement lines up definite possibilities for Da-sein, since attunement always pushes Da-sein in some direction. Da-sein’s possibility is still entrusted to itself.

Da-sein is in how it understands to be, through which it knows its possibilities. It knows these not through self-perception but because understanding is the there whose being has this knowledge. Understanding is essentially attuned, surrendered to thrownness. In understanding, Da-sein is its there and thus loses sight of itself for its there, so this thrownness means Da-sein has always already failed to recognize itself, and thus is always faced with the possibility of finding itself again.

The totality of relevance reveals itself as possible connections of things at hand, and even nature (as the “unity” of objective presence) is only discoverable as one of its possibilities discloses itself. Why does understanding happen in terms of possibilities? It has the existential structure of project, the “existential constitution of being in the realm of factical potentiality of being.” Da-sein, as thrown, always has and will understand itself in terms of possibilities as long as it is. This does not mean that understanding thematically grasps the possibilities themselves—that makes possibility a mere given. Project lets possibility be possibility, and as this, understanding is the mode of being of Da-sein in which it is its possibilities as possiblities.

This kind of being means Da-sein is always “more” than it actually is, rendering attempts to “catch” it as objectively present problematic. It’s never more than it factically is because it factically exists as its possibilities.

The being of the there gets its constitution through understanding and prject, is what it does or does not become, and so it makes sense for it to become what it already is.

Da-sein can understand itself as for-the-sake-of-which, as itself, authentically, or it can understand itself inauthentically, in terms of the world. This does not mean Da-sein is cut off from itself, since “world belongs to its being a self as being-in-the-world.”

Understanding works as an existential meaning of sight—not perception with eyes or nonsensory awareness of objectively present things, but allowing accessible beings to be encountered in themselves without being concealed. Every sense does this in its own way. Sight related to existence is transparency, not focus on a central self but simultaneous disclosure of being-in-the-world in all its constitutive factors.

Attuned understanding. Thrown project. We must develop these existentials.

0.32 Understanding and Interpretation

Interpretation comes existentially out of understanding, developing the possibilities there projected. We shall now pursue the phenomenon of interpretation in the understanding of the world. In taking care of things, Da-sein learns to understand them in their relevance. Thus out of understanding the world in taking care of it Da-sein interprets it. Things at hand in circumspection are interpreted in their in-order-to, and taken care of accordingly. What has been thus explicitly understood has the structure of something as something, meaning that circumspectly the thing as hand is for …, not just as a name but to understand it as that which is for . . .. To circumspectly “see” this thing as a table, door, etc. is not to interpret a statement, but a seeing of what is at hand, a process that is already understanding and interpretative. When we fail to understand something, fail to see the “as”, it is not more primordial than the simple seeing, which understands, but a derivation from it, akin to an obstacle. The “as” is the a priori existential constitution of understanding.

Again we must clarify that interpretation by no means merely “tacks on” or “adds” meaning to some objectively present “stuff”—the relevance disclosed in understanding the world and made explicit through understanding is there always already.

This “making explicit” means that, since we always already understand things at hand in terms of a totality of relevance, everyday, circumspect interpretation is based on that fore-having, grounded in foresight that works toward what has already been understood.

What is the structure of the “fore”? How is it related to the “as”?

When Da-sein discovers innerworldly beings, understands them, we say they have meaning, since they “can be articulated in disclosure that understands”. Meaning is structured by fore-having, fore-sight, and fore-conception, and something becomes intelligible as something in terms of meaning. Meaning is an existential of Da-sein, is exclusive to Da-sein, is the fulfillment of being-in-the-world through the beings discoverable in the world. Only Da-sein can be meaningful or meaningless. This is, again, not a value judgement.

Now this business of fore-conception is distressing, since it suggests a circularity in understanding. The circle of understanding is the expression of the existential fore-structure of Da-sein, and we thus must be aware of the primordial knowledge therein. Fore-having, fore-sight, and fore-conception—these are not to be established carelessly and arbitrarily but through working in terms of the things themselves. Understanding is existentially the potentiality for being of Da-sein, and this is not the place for rigorous exactness!

0.33 Statement as a Derivative Mode of Interpretation

Statement (which we have encountered already in its modification of the “as”) is profoundly connected with the ontological problematic, as any look into the role of truth as logos in ancient ontology will attest. These three significations of the term statement are interconnected and comprise its full structure.

1. Pointing out, as in the primordial meaning of logos as apophansis—letting beings be seen from themselves. Statements not as meaning containers but for accessing beings in their modes of beig (even such as handiness). Pointing out always designates beings themselves rather than mere representations.

2. Predication—statement about a “subject”. What is stated here in “the hammer is too heavy” is the “hammer itself”, not in the subject but stated by the predicate, by the “too heavy”. This is a narrowing of pointing out, predications “are” as they do their pointing out. Subjects are determined by predicates. This determination limits seeing to what shows itself, as a mode of pointing out. When it faces something already manifest it dims beings down and lets the beings be seen in their definite character (the hammer’s heaviness).

3. Communication, letting someone see with us what has been pointed out and predicated. What is “shared” is the being toward what is pointed out, a being-in-the-world where the world is that from which what is pointed out is encountered. Further retelling can change what is pointed out as the statement’s being toward is not to the original being talked about but instead what is heard.

In our discussion of meaning we will ignore the confused and arbitrary phenomenon of “validity” and make to attempt to restrict meaning’s conception to “judgement”, but we will understand it as that which makes visible “the formal framework of what can be disclosed in understanding and articulated in interpretation.”

So what is the unified phenomenon these three significations comprise? “Statement is a pointing out which communicates and defines.” Now for the statement to be a mode of interpretation, some things have to hold true. We see that statements on their own cannot disclose beings in general, but always work from a being-in-the-world. What is pointed out in determining works from something already disclosed. Determining begins with directedness, inheriting the direction in which beings are envisaged, predicates are loosened. Just like interpretatio, the statement begins with fore-having, fore-sight, and fore-conception. But it is a derivative mode of interpretation. When a hammer is too heavy one might state it any different way, but every statement operates under the presupposition of a more primordial interpretaion that is inherent in circumspectly and heedfully exchanging the hammer for a more appropriate one. How does the statement originate from this? With the hammer, something at hand for … becomes that “about which” the statement is made. The being is held in fore-having, but fore-sight looks to objective presence. This objective presence covers over handiness in the statement and the “as” moves from reaching out into circumspection and the surrounding world, only seeing what is objectively present and what objectively present things can be determined from it. The statement turns the “as” of circumspect interpretation into the determination of objective presence. The former “as” is existential-hermeneutical and the latter is apophantical. But every statement has its origin in circumspect interpretation, so attempts to reduce them to theoretical propositional statements distort their meanings. Aristotle conceived logos as synthesis and diairesis, binding and separating. This is the phenomenon of “something as something”, where the somethings ar eunderstood together and taken apart. Synthesis and diairesis are “relating”. This shows us how the “logic” of logos is rooted in the existential analytic of Da-sein.

0.34 Da-sein and Discourse: Language

Existentially equiprimordial with attunement and understanding is discourse. What is articulated in speech is meaning, and what is articulated in discourse is the totality of significations. Discourse is the articulation of the intelligibility of the there. This connection to disclosedness grounds discourse in the world, gives it a worldly mode of being. Discourse is an expression of the attuned intelligibility of being-in-the-world. “Words accrue to significations. but word-things are not provided with significations.” Discourse is expressed in language, a totality of words. Through language discourse can be found as an innerworldly being.

Discourse maintains itself in a particular way of heedful being-with-one-another, talking in assenting, warning, interceding, etc. It is discourse about …, it communicates. Communication is not a giving or sharing of experiences, opinions, or wishes from inside one being to inside another. Communication is a being-with that already exists, being made explicit, grasped, and appropriated through discourse. All communicating discourse expresses itself, not from the inside as though cut off intially and isolated from the outside but because as being-in-the-world it is already “outside” as understanding. That is the “itself” being expressed: attunement, being outside, the there. Attunement is made known poetically, through “the way of speaking”.

What discourse is about, what is said, communication, and making known. These are not properties taken from language but rather they are existential characteristics of Da-sein that make language ontologically possible. Thus we cannot grasp “the essence of language”, cannot arrive at a definition of language by putting these together as though they are pieces of the whole—we must first develop the structure of discourse according to the analytic of Da-sein.

Hearing is an existential possibility of discourse. “Listening to … is the existential being-open of Da-sein as being-with for the other.” “Da-sein hears because it understands.” Da-sein is being-in-the-world that understands, along with Mitda-sein, and so listens and belongs to itself and others, listening to each other. This listening to each other is the grounding for being-with, and has the possibilities for following, agreeing, not hearing, defying, etc. What we hear are things, not bundles of “simple ideas’, a chaotic swirl of pure sensation. We hear a wagon or a motorcycle, and in discouse we initially understand what is said in the same way! Even if we don’t understand it, what we hear is understood as unintelligible words. Da-sein, more primordially than hearing, hearkens. Understanding doesn’t come from talking or listening, listening comes from understanding and being-with! Talking about something too much covers up understanding and gives false sensations of clarity. Keeping silent “lets something be understood” in an authentic way, so long as it is the reticence of someone who has something to say, so long as it happens in discourse. It allows for a genuinely transparent being-with-one-another.

The Greek emphasis on zoon logon echon arises from the fact that Da-sein as discoursing being-in has already expressed itself—it is a statement of language’s ontological role. Understanding it as “rational living being” ignores the phenomenal basis. Grammar and categories look to the statement and objective presence. Philosophy needs to look at language in its ontological character.

Now we will use all this and more to find a more ontologically primordial view of the everydayness of Da-sein.

B: The Everyday Being of the There and the Falling Prey of Da-sein—We will find thrownness through everydayness, which is found in Da-sein’s possibilities as the they.

0.35 Idle Talk

Time to ask about the mode of being of discourse that expresses itself. What is its being, and what does that tell us about everydayness? Discourse expressing itself is communication. In this language that expresses oneself there is an average intelligibility through which a listener can understand the discourse communicated without coming to a being toward what is being talked about, can achieve a primordial understanding, approximate and superficial. Listener and speaker mean the same thing because they understand what is said in the same averageness.

Since hearing and understanding happen befor ewhat is spoken about, before communication, the beingsaid, the dictum, guarantees the genuineness of the understanding belonging to discourse. In an average understanding discourse communicates through gossip and passing the word along, and becomes more authoritative the wider it is spread. This is idle talk. The average understanding neither wants nor needs distinction between mere gossip and grounded discourse, since it understands everything. It spares us the job of genuine understanding and develops an indifferent intelligibiliti that leaves nothing incomprehensible. Idle talk is not deliberate deceit, it does not have the kind of being of consciously passing off something as something else. Idle talk is a closing off, it omits going back to the foundation of what is being talked about. Worse, it thinks it has gone back, and discourages further investigation. Da-sein does this all the time, is always working in, from, and against everyday understanding. Public interpretation, the they, always decides the possibilities of Da-sein’s attunement. The “obviousness and self-assurance” of the average way of being interpreted conceals Da-sein’s plummet into groundlessness from itself.

0.36 Curiosity

Da-sein has a tendency toward “seeing”, toward genuinely appropriating beings in relation to its possibilities of being, called curiosity, which happens in an everyday way. In rest, circumspection turns away from its work and rout eof procedure and becomes free. Care becomes free circumspection which has nothing to bring near, and it makes what is nearest at hand suddenly distant, near only in outward appearance. Free curiosity does not see to understand; it sees merely to see. It leaps around, not staying with what is nearest but moving from novelty to novelty, maintaining the possibility of distraction. Curiosity never dwells anywhere, is everywhere and nowhere. When in the mode of being-in-the-world of curiosity, everyday Da-sein consantly uproots itself. Discourse and sight, idle talk and curiosity, understand everything and see everything in a supposedly genuine way. When this genuineness is supposed, a third phenomenon appears to disclose everyday Da-sein.

0.37 Ambiguity

Suddenly there is ambiguity as to what is genuine understanding and what is not. Also, not only does everyone already understand what is the case but they also know what needs to be done and what others think and feel. Idle talk is indignant toward “what needs to be done” actually happening; everyone knows they could have done the same, since they figured it out, and when something is done there is no more opportunity for guessing. Idle talk is so quick and new in its distractedness anyway that it leaves behind any actual doing. Ambiguity passes talking and guessing off as what is happening and calls taking action something unimportant which comes after. Idle talk intrudes into being-with-one-another and causes us to keep tabs on one another. This comes not from Da-sein but from the thrownness of being-with-one-another.

The horizon of all these structures reveals to us the basic kind of being of everydayness.

0.38 Falling Prey and Thrownness

Idle talk, curiosity, and ambiguity reveal a basic kind of the being of everydayness called the entanglement of Da-sein, meaning that Da-sein is together with the “world” it takes care of. Da-sein has the possibility of being itself, meaning that it has initially fallen prey to the “world” and become lost in the publicness of the they. Thus idle talk, curiosity, and ambiguity guide us into an inauthentic being-with-one-another. “Inauthentic” does not mean that Da-sein totally loses its being in everyday being, it is simple a kind of being-in-the-world which is totally taken in by the world, Mitda-sein, and the they. Not-being-its-self is a positive possibility of beings absorbed in and taking care of a world. This nonbeing is actually the kind of being nearest to and most common for Da-sein. Entanglement is not a “fall” from a purer and higher “primordial condition”. Da-sein, in falling prey, falls away from itself, has already done so in its factical being-in-the-world, and falls prety to the world belonging to its being. This is an existential determination of Da-sein itself, not a statement about objectively present relations or some horrible ontic quality society ought to move beyond.

Being-in-the-world presents Da-sein with the possibility in idle talk and public interpretedness of losing itself in the they, with the constant temptation of falling prey. Bein-in-the-world is tempting. Idle talk as having-seen-everything and ambiguity as having-understood-everything convince Da-sein that it knows the full breadth of possibilities of its being and that it is leading a full and genuine “life”. By convincing Da-sein that everything is in “the best order”, entangled being-in-the-world is tranquilizing. As it “understands” everything, Da-sein compares itself with everything and thus conceals its ownmost potentiality for being-in-the-world. Entangled being-in-the-world is also alienating!

Rather than driving Da-sein from itself, this alienation drives Da-sein into itself in a “self-dissection” that explores countless possibilities of interpretation, entangling it in itself. Da-sein plunges out of itself into itself, into inauthentic everydayness. Understanding is constantly town away from authenticity into the they, still pretending to be authentic. In fallig prey, Da-sein in eddying.

Da-sein remains in the throw, is sucked into the eddy of the they’s inauthenticity, because of its facticity. Thrownness belongs to Da-sein, which is always concerned in its being about that being. This only contradicts itself if Da-sein is an I-subject separate from the world, which it is not. “Da-sein can fall prety only because it is concerned with understanding, attuned being-in-the-world.” Authentic existence is existentially just a modified form of everydayness.

Now we pick apart care with the same infuriating depth and intensity.

Care as the Being of Da-sein

0.39 The Question of the Primordial Totality of the Structural Whole of Da-sein

Being-in-the-world is whole. How is the totality of this whole to be determined existentially? Da-sein is disclosed to itself in thrownness, its mode in which it always is its possibilities. Being-in-the-world is always for itself, but this self is the inauthentic they-self, so being-in-the-world is always entangled. Da-sein is entangled-disclosed, concerned with its ownmost potentiality of being together with the “world” and in being-with with the others.

But we won’t get what we want by “cobbling together elements”. We must pursue it according to that understanding attunement in which Da-sein is disclosed to itself in the most far-reaching and most primordial way—Angst. This will be worked from as a form of entanglement rather than fear. “Its being reveals itself as care.” We now have to demonstrate that as soon as Da-sein expresses anything about itself it has already interpreted itself pre-ontologically as care. Then we go all over the fucking place, and I’ll cross theose bridges when I come to them. Seriously, JESUS.

0.40 The Fundamental Attunement of Angst as an Eminent Disclosedness of Da-sein

How is Da-sein brought before itself in Angst through its own being in a way that phenomenologically defines the being there disclosed in its being?

Da-sein can flee from itself because it is brought before itself in the first place. So while existentielly entanglement leads away from Da-sein, existentially and ontologically there is a turning toward what Da0sein flees in a phenomenologically interpretive way. In fact, this fleeing ensures that when orienting our analysis toward the phenomenon of entanglement we will not find an artificial self-conception of Da-sein. So all this has some merit, and it’s time to really look at Angst.

“In falling prey, Da-sein turns away from itself.” This shrinking back isn’t fear—it shrinks back from itself, and what is “fearsome” is always an innerworldly being. Da-sein is threatened by a being with the same kind of being as itself, a turning away based on Angst which makes fear possible. Once has Angst about being-in-the-world itself. In Angst, “the totality of relevance discovered within the world of things at hand and objectively present” is insignificant, is not relevant to the threat. There is no “over there” for what is threatening because it is nowhere, comes from no direction within nearness, is already “there” and yet nowhere at all. What is innerworldly is so insignificant that the only worldly thing left is the world itself, and that is what Angst is about. The world itself, “the possibility of things at hand in general.” What is ontically “nothing at all” is “none of the things at hand”, or rather the primordial “something” of the world. The world ontologically belongs to the being of Da-sein as being-in-the-world. “That about which Angst is anxious is being-in-the-world itself.” Angst, as not-having things at hand or Mitda-sein, removes the possibility of understanding, of falling prey. Da-sein is thrown back on its authentic potentiality-for-being-in-the-world. Being-with …, being familiar with …, and the tranquilized self-assurance in the everyday publicness of the they are a “being-at-home” standing in contradistinction to the uncanniness felt in Angst. “Being-in enters the existential ‘mode’ of not-being-at-home.” The flight of falling prey is toward innerworldly beings where “taking care of things, lost in the they, can linger in tranquillized familiarity”. Angst is essential to being-in-the-world, and is thus more primordial than everydayness. Familiar being-in-the-world is a mode of the uncanniness of Da-sein.

The unique individualizing of Angst fetches Da-sein out of falling prey and reveals authenticity and inauthenticity as possibilities of its being.

0.41 The Being of Da-sein as Care

The complete phenomenon of Angst, characterized equiprimordially by existentiality, facticity, and falling prey, constitutes the totality of the structural whole of Da-sein. As Da-sein looks to its potentiality-for-being, it is “beyond itself”, it is in a state of being-ahead-of-itself. This isn’t something done by a worldless “subject”—it is being-ahead-of-itself-in-already-being-in-a-world. The referential totality of the in-order-to isn’t an objectively present world attached to a subject but the phenomenal expression of Da-sein’s primordial wholeness. “Existentiality is essentially determined by facticity.” In being-ahead-of-oneself-already-being-in-a-the-world, one is essentially already in entangled being-together-with innerworldly handy things taken care of. Here’s our structure!

Being-ahead-of-oneself-already in as being-together-with. This demonstrates the full significance of care, used ontologically. Da-sein has the possibility of being free for athentic existentiell possibilities, a freedom which allows Da-sein to be related to its possibilities unwillingly and inauthentically, being-ahead-of-itself where the “self” is the they-self. Since Da-sein’s potentiality-for-being has the mode of being of being-in-the-world it is there that the relation to innerworldly beings is. Willing grasps somehing understood as something to be taken care of or “brought to its being through concern.” Something willed comes out of a for-the-sake-of-which, and through willing we can see the underlying totality of care showing.

Da-sein’s projecting is together with a world, where from it takes its possibilities. Average trnquillized everydayness levels down these possibilities, spurring busyness to create the illusion that something is happening when in fact no new possibilities are being willed. This tranquillized “willing” is a modification of being toward one’s potentiality-for-being where being toward possibilities becomes an act of wishing, where possibilities are ungrasped and unexpected. In wishing what is available is never enough, and is lost sight of. Wishing hankers after possibilities, an act which simultaneously closes off factical possibilities. What is “there” in wishful hankering, what is “never enough” becomes the “real world”. Wishing, like willing, presupposes care. Hankering modifies being-ahead-of-itself-in-already-being-in into being-in-he-world-already-among-…, revealing Da-sein’s predilection to be “lived” by the world in which it actually is. In predilection one is drawn in by what is hankered after. The urge “to live” is “toward at any cost”. It crowds out other possibilities. Both urge and predilection are rooted in thrownness, and can’t be removed. They are also ontologically based in care, which ontically modifies them as something authentic.

Now to show that this ontologically “new” concept of care is not contrived, but is actually very old and familiar.

0.42 Confirmation of the Existential Interpretation of Da-sein as Care in Terms of the Pre-ontological Self-interpretation of Da-sein

Uh... citations and reiterations. Nothing interesting. Wow. I guess mostly h shows old understandings of the term “cura” which demonstrate old pre-ontological musings on care as the stuff of human lives. Or they heal around a thousand hit points.

0.43 Da-sein, Worldliness, and Reality

We’ve touched before on objective presence as smoething that is true, is encountered, and is a part of Dasein’s being and being in general. It sometimes starts to sound like an alternate take on or misunderstanding of being, but objectively present reality is not “one kind of being among others”, but something that “stands ontologically in a definite foundational context with Da-sein, world, and handiness.” It’s time to tackle the problem of reality. We do this by examining

a: Reality as a Problem of Being and the Demonstrability of the “External World”

b: Reality as an Ontological Problem

c: Reality and Care

with regard to the question of fundamental ontology.

A—What does reality signify? This has often been bundled with “the external world”, sothat what is looked to is that which is “independent” or “external” to our consciousness. Still, the means of grasping it is said to be intuitive cognition, an act of the consciousness. To grasp transcendence we must grap the “independence” from consciousness of the real as well as to what extent this consciousness is itself clarified with regard to its being. Then we may ontologically grasp the kind of being of transcendence and whether cognition is up to the task of accessing what is real.

The nature of Da-sein throws a gargantuan spanner in the works of this inquiry. Cognition is founded upon innerworldly beings; they are the means of accessing reality, and such access happens based ontologically on being-in-the-world, which is care. Da-sein has already accessed the world in its being! Objective presence can be covered over, but it is still only discoverable on the basis of the world already disclosed, is only even concealable on this basis! We are in a spot. Kant complains about how we can’t prove the existence of these things outside of us, and then... Well, now we’re getting into it again.

Time indicates something objectively present that persists, but it’s not me because my own “existence” (Kant uses the term for objective presence) in time is detrmined according to it. What it is is the possibility of te objectively present change “in me”. “The experience of the being-in-time of representations equiprimordially posits changing things ‘in me’ and persisting things ‘outside of me’.” Of course, the very language used stinks like shit.

Ah! And here he says it at last: “The objective presence together of the physical and the psychichal is ontically and ontologically completely different from the phenomenon of being-in-the-world.” Fuck, dude, you can ignore everything before that sentence. Proofs such as Kant desired miss the point, as usual. Da-sein already is in its being what Kant is trying to prove for it, and what Kant is trying to work from is not up to the task. Appeals to faith are still trying to satisfy demand for such proofs, thus maintaining all the faulty presuppositions.

Being-in-the-world does not deny or refute the real objective presence of the external world, but it does deny the need or possibility for a proof of the “world’s” reality. Oh, thank CHRIST. We get to go back into the existential analytic, back into ontology. Fuck you for this, Marty.

B—Dilthey describes reality as resistance, something experienced in impulse and will. Impuls and will are out to get … and they encounter resistance, which means this is together with a totality of relevance.

“…[T]he discovery of resistance in striving, is ontologically possible only on the basis of the disclosedness of world.” Furthermore, this striving is a modification of care, not something that “emerges” on its own. And so, uh, FUCK DESCARTES.

C—Still not saying shit. Time to cut rope from this shitty section.

0.44 Da-sein, Disclosedness, and Truth

Truth! It’s wound up with being and Da-sein, with Da-sein’s understanding of being.

a: The Traditional Concept of Truth and its Ontological Foundations—Starting with Aristotle, truth was accepted as the agreement of a proposition (judgement) with its object. Well, every “agreement” is a relation, yet we are trying to relate something real with something ideal, a problem by now too familiar. Really what is demonstrated is the being-discovered of the being itself, “that being in the how of its being discovered.” The being itself is what is stated.

A true statement discovers the beings in themselves, letting them “be seen” in their discoveredness. A statement’s truth is understood as discovering. This is not agreement as with correspondence of one being to another. Hey, guess what! This all depends on being-in-the-world!

b: The Primordial Phenomenon of Truth and the Derivative Character of the Traditional Concept of Truth—Truth means to-be-discovering, unconcealment. The phenomenon of truth is witnessed most primordially in the ontological foundations of discovering. Da-sein is discovering, and innerworldly beings taken care of are discovered. Only through the disclosedness of Da-sein is the most primordial phenomenon of truth attained. Da-sein is “true”, is “in the truth” in that it is its disclosedness. This truth isn’t some allencompassing, omniscient Logos in the sky—it simply means that “the disclosedness of its ownmost being belongs to its existential constitution.” What this means is:

1. Disclosedness in general is essential to the constitution of Da-sein’s being. The phenomenon of care has made explicit the totality of the structure of being. Care belongs to both being-in-the-world and being together with innerworldly beings.

2. Thrownness reveals that Da-sein is in a definite world “together with a definite range of definite innerworldly beings. Disclosedness is essentially factical.”

3. Project is disclosing being toward Da-sein’s own potentiality-of-being. Da-sein understands itself in terms of the “world”, others, or in terms of this potentiality. The latter possibility means Da-sein discloses itself to itself in and as this potentiality. This is an authentic disclosedness of primordial truth in authenticity, and the most primordial and authentic disclosedness in which Da-sein can be in this way is the truth of existence. So basically it’s only through analysis of Da-sein’s authenticity that it becomes ontologically and existentially definite.

4. Falling prey is Da-sein losing itself in its “world”. Public interpretedness, in the form of idle talk, curiosity, and ambiguity, closes off what is disclosed and disguises what is discovered. Beings show themselves, but as illusions. Here Da-sein is in “untruth” in accordance with the constitution of its being. These terms are used ontolotgically, with no ontically negative value judgement. When the statement “Da-sein is in the truth” is fully developed ontologically it also says “Da-sein is in untruth”. Only insofar as innerworldly beings are discovered already with Da-sein are they covered over and given the possibility of being encountered.

Discovering a being’s truth robs it of its concealment, of its being-in-untruth. We can now say that truth, most primordially, is the disclosedness of Da-sein, which has the discoveredness of innerworldly beings. Also, Da-sein is equiprimordially in truth and untruth.

As for the truth of statements, they ar ebased on understanding and the disclosedness of Da-sein, and so their truth is accordingly based on the disclosedness of understanding. Innerworldly beings are discovered through care, and Da-sein expressed itself as discovering. It does this expressing in statements, which “communicate beings in the how of their discoveredness”. Statements preserve discoveredness in what is expressed, which itself becomes an innerworldly thing at hand. Because of the preservation, even repetitions bring Da-sein into a being toward the beings originally discussed, though Da-sein does not repeat the act of discovering.

So discoveredness can be appropriated without “original” experience through hearsay. In this case, the relation of statement to the beings it discovers is presented as an objecively present thing relating objectively present things. This is how the traditional concept of truth as objectively present agreement of statement to object is brought about. This is its ontological derivation.

“The statement is not the primary ‘locus’ of truth, but the other way around.”

c: The Kind of Being of Truth and the Presupposition of Truth—Since truth comes out of Da-sein’s discoveredness, “all truth is relative to the being of Da-sein.” Da-sein presupposes in being ahead of itself, and so the truth presupposed has the kind of being of Da-sein. Discoveredness is presupposed in Da-sein’s project.

Skeptics refute themselves insofar as they are at all, and udnerstand themselves in their being. Skepticism is refuted out of groundlessness, not formal dialectic. The a priori is not the eternally true, but the equiprimordially true and untrue.

Has this shit worked at all? Find out next time!

Da-sein and Temporality

0.45 The Result of the Preparatory Fundamental Analysis of Dasein and the ask of a Primordial, Existential Interpretation of This Being

The being of Da-sein is contained in care. Existence means Da-sein is as an understanding potentiality-of-being, concerned in its being about its being. Care helped develop existence’s equiprimordial connection with facticity and falling prey. But here we have left out authentic potentiality-of-being, and thus cannot claim that our existential analytic is primordial. We need Da-sein as a whole, we must ask the question of its potentiality-for-being-a-whole. The potentiality is always outstanding as long as Da-sein is. Belonging to the outstanding is the “end” itself, death. In being a whole, Da-sein is in the mode of being-at-an-end.

How can Da-sein exist as an authentic whole? Conscience? Wanting-to-have-a-conscience/ Using the word VERY FUCKING CAREFULLY we can see an authentic potentiality-of-being here, and it’s made definite existentielle by being toward death.

Da-sein comes out of temporality, which it encounters through care as “time”.

Basically we take Da-sein out to being in the horizon of time. OK, let’s do it!!

The Possible Being-a-Whole of Da-sein and Being-toward-Death

0.46 The Seeming Impossibility of Ontologically Grasping and Determining Da-sein as a Whole

Even when Da-sein has “settled its accounts” its being is influenced by “being ahead of itself”. Da-sein essentially has in its constitution a constant unfinished quality. As lon as it is, Da-sein has something outstanding, so that if it ever achieves “wholeness” it loses being-in-the-world, and is no longer experienced as a being. Ontological wholeness of Da-sein is not even graspable. We have to existentially develop the being-toward-the-end of Da-sein and death itself.

0.47 The Possibility of Experiencing the Death of Others and the Possibility of Grasping Da-sein as a Whole

Da-sein can neither experience nor understand as experienced its own death. Da-sein can experience another’s death through being-with wth others. What is experienced is a being’s transition from the kind of being of Da-sein to no-longer-being-there, ending being qua Da-sein and beginning being qua objective presence. Note that it’s not merely lifeless material, but an unliving thing no longer alive.

But in a funeral we are with someone not factically “there”. We can only be with him in terms of this world. The deceased has left this world, but we cannot experience that. “We are asking about the ontological meaning of the dying of the person who dies, as a potentiality-of-being of his being, and not about the way of being-with and the still-being-there of the deceased with those left behind.”

Death is an existential phenomenon. Going-out-of-the-world of Da-sein is dying, that of what is only alive is perishing.

0.48 What is Outstanding, End, and Totality

The story so far:

1. As long as Da-sein is it has a not-yet, what is constantly outstanding.

2. “The coming-to-its-end of what is not-yet-at-an-end…has the character of no-longer-being-there.”

3. Coming-to-an-end is a mode of being in which Da-sein cannot be represented by another.

What is outstanding? It “belongs” to Da-sein, but lacks. It’s not yet available. This not-yet is “filled out until the sum owed is ‘all together.’ ” Outstandingness means what belongs together isn’t. “The being for which something is outstanding has the kind of being of something at hand.” The together or the untogether based on it, is characterized as a sum. This can’t define the not-yet that is Da-sein’s own possible death since it has the kind of being of something at hand. The together of Da-sein’s being “in running its course” isn’t a gradual piecing-on of beings at hand.

“In order to…define the being of the not-yet of the character of Da-sein, we must reflect on beings to whose kind of being becoming belongs.” Unripe fruit becomes ripe, it doesn’t piece itself together as something not-yet-objectively-present. The not-yet of unripeness isn’t something objectively present that sits indifferently with the fruit, but the fruit itself in a kind of being. In this way, Da-sein is always already its not-yet as long as it is.

But death is not like ripeness—it does not imply fulfillment, or even an ending. Fulfillment is founded upon “finishedness”, which is possible only as adetermination of something objectively present or at hand. But death is nothing like the end of rain, or of a sandwich. Da-sein is not objectively present or at hand!

Da-sein already is its end. Da-sein is not always being-at-an-end, but being toward the end. At birth Da-sein is old enough to die—that is, as soon as it is Da-sein has death as a possible way to be.

So we haven’t tied this together yet in terms of the kind of being of Da-sein. That means we must follow the phenomenon of care.

0.49 How the Existential Analysis of Death Differs from Other Possible Interpretations of This Phenomenon

How death is defined in terms of life is always ontically understood, and we must look at this understanding from the ontology of Da-sein. Da-ein, as Da-sein, never simply perishes, since it is not “merely” alive. Its physiological death is its demise, which can only happen through dying (being toward death).

“The existential analysis [of death] is methodically prior to the questions of a biology, psychology theodicy, and theology of death.” We have to ontologically characterize the kind of being in which the “end” enters average everydayness. Now we must ontologically look at Da-sein’s being-toward-the-end.

0.50 A Preliminary Sketch of the Existential and Ontological Structure of Death

Time to define death in terms of the characteristics of its being:

in being-ahead-of-itself , existence

in already-being-in …, facticity

in being-together-with …, falling prey

How are the existence, facticity, and falling prey of Da-sein revealed in the phenomenon of death? The not-yet was inappropriate, since the most extreme not-yet has the character of something Da-sein relates to. Death isn’t not yet objectively present, death is an imminence. Other things can be imminent, from objectively present or at hand beings to Mitda-sein to possibilities-of-being, but death as the possibility of no-longer-being-able-to-be-there throws Da-sein back on its ownmost potentiality of being. Death makes Da-sein imminent to itself, and all relations to other Da-sein disappear. Da-sein cannot evade the possibility of death. It is an eminent imminence, Da-sein’s ownmost nonrelational possibility not to be bypassed. Beingtoward-death is the most primordial concretion of care, of being-ahead-of-itself. Death is a possibility Da-sein is thrown into without knowledge, so death belongs to being-in-the-world. This thrownness into death reveals itself to Da-sein in the attunement of Angst, Angst about death isn’t dear of one’s demise but the disclosedness of Da-sein’s existence as thrown being-toward-its-end. So the existential concept of dying is thrown being toward the ownmost potentiality-of-being not to be bypassed. The flight from being-toward-death makes itself known in the entangled being together with … of factical existing. Ontologically, dying is grounded in care. Now we have to show the connection in everydayness.

0.51 Being-toward-Death and the Everydayness of Da-sein

How does the attuned understanding in the idle talk of the they have disclosed being-toward-death? Public, everyday being-with-one-another understands death as a familiar event in the world, keeping it inconspicuous. The idle talk that understands it thusly says: “One also dies at the end, but for now one is not involved.” “One dies” is clearly of the they, as it makes death something “all around us”, something happening to the they but never I myself. Death is no longer mind, but it “concerns” me. This is ambiguous in the way of idle talk.

While a person is dying, the others in his they (“neighbors”) try to convince him he will evade death, making sure of a constant tranquillization toward death for him and for them. The they also distorts Angst into fear of a future event, a fear it encourages Da-sein to ignore and forget.

This all has the kind of being of falling prey. Everyday being-toward-death is a flight from death. “Being toward the end has the mode of evading that end—reinterpreting it, understanding it inauthentically, and veiling it.” But all that is lost here is Angst. Da-sein is still concerned with its death.

How does this evasive Da-sein understand its own death?

0.52 Everyday Being-toward-Death and the Complete Existential Concept of Death

Everydayness ambiguously acknowledges death as both “certain” and “also sometime, but for the time being not yet.” What is this certainty? Certainty is being-certain as a kind of being of Da-sein, and beings Da-sein is certain of are also “certain”. In conviction, Da-sein understands the thing discovered certainly according to its own testimony. Certainty can keep Da-sein in truth if it is based on the discovered beings themselves. The covering over of death isn’t undertainty in the sense of doubting, but in the sense that is is a certainty inappropriate for holding-for-true. Inappropriate certainty covers over that of which it’s certain. The certainty of death as an event in the surrounding world is inappropriate as a way to get at being-toward-the-end, so we see how everydayness covers over death and makes it “certain” at the same time.

But when they say death is certain, Da-sein is under the illusion that it is itself certain of its own death, even if Da-sein is not even certain of its ownmost nonrelational potentiality-of-being not-to-be-bypassed. Entangled everyday Da-sein “knows about the certainty of death, yet avoids being-certain.” Everydayness also says “but not yet”, covering over the fact that death is possible in every moment. Everyday beingtoward-death says the when of death is definite, trying to evate its indefiniteness. it does not make the when definite by saying wen the demise will occur but by putting before death in Da-sein’s attention “those manageable urgencies and possibilities of the everyday matters nearest to us.” This covering-over of death’s indefiniteness also covers over its certainty, thus covering over the ownmost character of the possibility of death.

So now we can define the full existential and ontological concept of death:

As the end of Da-sein, death is the ownmost nonrelational, certain, and, as such, indefinite and not to be bypassed possibility of Da-sein.

Da-sein exists toward its death, and this most extreme not-yet is always already included in Da-sein; “The phenomenon of the not-yet has been taken from the ahead-of-itself.” Now we turn our attention to the ontological possibility of authentic being-toward-death.

0.53 Existential Project of an Authentic Being-toward-Death

We have established the existential concept of death as well as inauthentic being-toward-death, so we know that to which authentic being-toward-death relates as well as that which it is not. Authentic being-toward-death can not evade death. Being-toward-death is a being toward a possibility, which can mean to be out for something possible, to take care of its actualization. Even when a possibility has been actualized it remains something possible for …, it is characterized by an in-order-to. Being out for something and taking care of it “circumspectly looks away from the possible to what it is possible for.” It is different with death, however, since rather than a possible thing at hand or ojectively present it is a possibility-of-being of Da-sein, and furthermore since taking care of the actualization of the possibility of death deprives Da-sein of its ground for existing being-toward-death.

Da-sein relates to what is possible in its possibility by expecting it. This is not just a looking away from what is possible to its possible actualization but a waiting for that actualization. One expects the reality of what is expected, and the possible is thus connected to the real in expecting.

But in being toward death as possibility, in anticipation of this possibility, we see that the possibility of death is as far removed as possible from anything real. The possibility of death is the possibility of the impossibility of exitence in general. Anticipation, relating Da-sein to its possibilities as possibilities, shows itself as the possibility of authentic existence. Thus we must set forth the concrete structure of anticipation of death, by defining the role of anticipatory disclosure in understanding oneself in the potentiality-of-being-at-an-end. Death individualizes Da-sein down to itself, not death but its nonrelational character understood through anticipation.

Da-sein is authentically itself when it projects itself as being-together with things taken care of and concernful being-with … upon its ownmost potentiality-of-being and not the possibility of the they-self.

Anticipation also does not evade the imminence of death but frees itself for it. This keeps Da-sein from “falling back behind itself”, or (Nietzsche) “becoming too old for its victories”. It prevents a finite understanding of Da-sein’s existence from occulting the influences of others’ existence-possibilities. The individualization of death as the nonrelational possibility not-to-be-bypassed discloses the potentialities-of-being of others, and anticipation of the possibility not-to-be-bypassed discloses all possibilities lying before it. This anticipation thus gives Da-sein the possibility of existing as a whole potentiality-of-being, of taking its whole in advance in an existentiell way.

Anticipation also does not evade the imminence of death but frees itself for it. This keeps Da-sein from “falling back behind itself”, or (Nietzsche) “becoming too old for its victories”. It prevents a finite understanding of Da-sein’s existence from occulting the influences of others’ existence-possibilities. The individualization of death as the nonrelational possibility not-to-be-bypassed discloses the potentialities-of-being of others, and anticipation of the possibility not-to-be-bypassed discloses all possibilities lying before it. This anticipation thus gives Da-sein the possibility of existing as a whole potentiality-of-being, of taking its whole in advance in an existentiell way.

Anticipation also does not evade the imminence of death but frees itself for it. This keeps Da-sein from “falling back behind itself”, or (Nietzsche) “becoming too old for its victories”. It prevents a finite understanding of Da-sein’s existence from occulting the influences of others’ existence-possibilities. The individualization of death as the nonrelational possibility not-to-be-bypassed discloses the potentialities-of-being of others, and anticipation of the possibility not-to-be-bypassed discloses all possibilities lying before it. This anticipation thus gives Da-sein the possibility of existing as a whole potentiality-of-being, of taking its whole in advance in an existentiell way.

How, existentially, is this constant threat disclosed? “The attunement which is able to hold open the constant and absolute threat to itself arising from the ownmost individualized being of Da-sein is Angst.” Angst is anxious about one’s potentiality-of-being, disclosing the most extreme possibility. Being-toward-death is essentially Angst.

Summary: “Anticipation reveals to Da-sein its lostness in the they-self, and brings it face to face with the possibility to be itself primarily unsupported by concern taking care of things, but to be itself in passionate anxious freedom toward death which is free of the illusions of the they, factical, and certain of itself.”

We now have the ontological possibility of an authentic potentiality-for-being-a-whole. We must make an essential connection with that authentic potentiality-of-being attested to.

The Attestation of Da-sein of an Authentic Potentiality-of-Being, and Resoluteness

0.54 The Problem of the Attestation of an Authentic Existentiell Possibility

We now know the character of what we are looking for, but it’s time to look for it. We want “an authentic potentiality-of-being of Da-sein that is attested by Da-sein itself in its existentiell possibility.” This is going to be rooted in the being of Da-sein. We look toward an authentic potentiality-of-being-one’s-self, which “self” is a way of existing, the who of Da-sein, which for the most part is not I myself but the they-self. Authentic being-a-self is an existentiell modification of the they. In being lost in the they, Da-sein’s choices are silently made for it, and not by any definite “who”. In this mode, when choices are made for it by no one in particular, Da-sein is caught up in inauthenticity. This is reversed by an existentiell modification whereby Da-sein chooses to make this choice, makes possible Da-sein’s authentic potentiality-of-being. First Da-sein must find itself to get out of lostness in the they, it must be “shown” to itself in its possible authenticity, a potentiality-for-being-its-self which Da-sein already is but has not yet had attested to it. What does this attesting is the “voice of conscience”, a phenomenon we shall investigate existentially and with the aim of fundamental ontology. Careful! This is not psychology, biology, theology, or any such dissolution of the phenomenon. This will be hyped. Conscience is a call which summons Da-sein to its ownmost potentialityof-being-a-self, by summoning it to its ownmost quality of being a lack.

0.55 The Existential and Ontological Foundations of Conscience

Conscience “gives one to understand something”. Da-sein is summoned by a call that calls silently, unambiguously, and with no foothold for curiosity. In other words, the call is the hearing that summons Da-sein from the listening to the they, and its character is in every way opposite such hearing. The call interrrupts, summoning Da-sein from idle talk. This calling is a mode of discourse.

0.56 The Character of Conscience as a Call

Discourse draws from what is talked about and makes it accessible to the Mitda-sein of others. What is talked about in the call of conscience, what is summoned? The they-self is summoned to the self, not to close itself off from the “external world” but to summon the self which is simply being-in-the-world. But what is what is spoken here? Conscience has nothing to tell and does not “converse with itself” in the self summoned.

Conscience speaks solely and constantly in the mode of silence.

What the call discloses is unequivocal, even though it is interpreted differently in individual Da-sein according to its possibilities of being understood. The content of the call is “seemingly indefinite” but its direction is sure. The call is only “deceptive” when it is not understood authentically and is instead drawn by the they-self into conversation with one’s self, distorting it in its character of disclosure. Conscience as call is a summons to the they-self in its self. It is thus a summons of the self to its potentiality-of-being-a-self, calling Da-sein forth to its possibilities.

0.57 Conscience as the Call of Care

“Conscience calls the self of Da-sein forth from its lostness in the they.. . .The call passes over what Da-sein understands as initially and for the most part in its interpretation in terms of taking care of things.” The caller of the call is never familiar, considered, or talked with or about. The call comes from Da-sein, over Da-sein. It is something totally unexpected that we are unprepared for. We need to “hold on” to the notion of conscience as a phenomenon of Da-sein, so that this being is the only way to interpret the kind of being of the “it” that calls. Da-sein flees from uncanniness and thrownness to attunement and the they-self.

The conscience is Da-sein, calling out from uncanniness. The voice is alien to the they-self, definable by nothing “worldly”. “It calls, and yet gives the heedfully curious ears nothing to hear that could be passed along and publicly spoken about.” This call individualizes Da-sein to itself in its uncanniness, making possible for Da-sein its project upon its ownmost potentiality-of-being.

Conscience is the call of care: Da-sein calls, anxious in thrownness about its potentiality-of-being, summoning Da-sein (itself), calling it forth to its ownmost potentiality-of-being. What the summons calls forth is Da-sein, out of falling prey to the they. Conscience has its ontological possibility in the fact that Da-sein is care in the ground of its being. Conscience is at bottom always mine.

We have it now as a phenomenon of Da-sein, but we still need it as an attestation in Da-sein of its ownmost potentiality-of-being. For this we must examing the hearing that goes along with the call. Then we can fins what the call gives to understand. There is also this curious connection with “guilt” to be explored.

0.58 Understanding the Summons, and Guilt

Guilt. We all hear it, are called “guilty”, but what is its primordial, existential meaning and how did that get distorted by everyday interpretation?

Guilt in the sense of “owing something” is related to things that can be taken care of. This and “being responsible for …” are “making oneself responsible”, by breaking a law. This can have the character of “becoming responsible to others”, and blah blah blah none of this is what we are after. Being resopnsible to others is being the ground for a lack in another’s Da-sein, a being-the-ground which is “lacking” in terms of that for which it is the ground. A lack is the not-being-objectively-present of something which ought to be and can be, according to taking care of things and the law. This is bound up with objective presence, so we must remove the idea of guilt from it.

While the idea of “guilty” is distanced from lacking, it still bears the quality of the not. So we may define our “formal existential idea of ‘guilty’ as being-the-ground for a being which is determined by a not—that is, being-the-ground of a nullity.” A lack is not a deficiency of a “cause”—“indebtedness is possible only ‘on the basis’ of a primordial being guilty.”

The being of Da-sein, care, includes facticity (thrownness), existence (project), and falling prey. Da-sein exists as thrown, and so can’t get away from its thrownness. This is not due to contingency, some“thrownness” that occurred as an event, but because Da-sein is its “that” as care. Da-sein did not lay the ground of its potentiality-of-being in which it rests. Da-sein projects itself upon the possibilities into which it is thrown, taking over being the ground in existing despite the fact that it can never overpower the ground. “Being its own thrown ground is the potentiality-of-being about which care is concerned.” As it is its ground, Da-sein is a nullity of itself.

When Da-sein stands in one possibility or another it is constantly not other possibilities. The project is thus essentially null in the sense belonging to Da-sein’s freedom for its existentiell possibilities. Care is so thoroughly permeated with nullity, is the ground of nullity, that we can say Da-sein as such is guilty.

The call of conscience is a summoning to being-guilty, a calling forth to the potentiality-of-being that I always already am as Da-sein.

The call calls Da-sein to listen to its ownmost possibility of existence—to choose itself. “Understanding the summons means: wanting to have a conscience.”

God DAMN I hope none of this ends up mattering.

So the “lack” of “guilt” in terms of objective presence turns into the potentiality-of-being that grounds Da-sein and yet which Da-sein can never be, it is the nullity of the project, of all the possibilities relinquished in the standing in this possibility.

0.59 The Existential Interpretation of Conscience and the Vulgar Interpretation of Conscience

Remember that even the vulgar interpretation of conscience is in touch with the phenomeon pre-ontologically. It’s not outright “wrong”, but simply keeping to what they know as conscience. This vulgar interpretation must thus be accessible from the ontological interpretation, so we must cover these four departures, where the vulgar interpretation disagrees with the idea of conscience as the summons of care to being-guilty:

1. Conscience has an essentially critical function.

2. Conscience always speaks relative to a definite deed that’s been done or wished for.

3. Experientially, the “voice” is never so radically related to the being of Da-sein.

4. What of “good” and “evil”? Doesn’t conscience “reprove” and “warn”?

Conscience is primarily “bad”, letting us know that “guilty” gets experienced first. This turns up after the realization or abandonment of the possibility of the deed. By calling back into thrownness, bad conscience points forward.

“The order of succession in which experiences run their course is not valid for the phenomenal structure of existing.”

“Good” conscience is interpreted as a privation of “bad” conscience,as a making certain that Da-sein “didn’t do it” and therefore is innocent. The “certainty” tranquilizes the desire to have a conscience, to understand one’s being-guilty. This is bogus. “Good” conscience is neither an independent nor a founded form of conscience—it is not a phenomenon of conscience at all. Here we can also distinguish between a forward-looking, warning conscience and a backward-looking, reprovig conscience. Both of these, however, are still there to deter Da-sein from a willed deed, and not in fact to summon at all.

The third complaint says that Da-sein’s everyday experience isn’t familiar with being summoned to be guilty. Still, vulgar conscience lacks ontological grounding, instead presupposing a moral law. This sees Da-sein as a being to be taken care of. The scope of eveyday conscience is inadequate for being the sole court for the interpretation of conscience.

So complaint two kind of falls apart here—while the call is frequently experienced as relating to objectively present “facts”, this understanding is restrictive and inadequate.

Complaint one speaks to an expectation that conscience will tell us what to do with a certainty of outcome. This would rob existence of the possibility of acting. The call is never “positive” or “negative” as something to be taken care of, since it concerns existence.

Anyway...

0.60 The Existential Structure of the Authentic Potentiality-of-Being Attested in Conscience

In understanding the call of conscience, in wanting to have a conscience, Da-sein is brough face to face with its uncanniness. Da-sein is ready for Angst. Recall how we said that in order to be silent one must first “hav something to say”. It’s this mode of discourse, reticence, in which the call brings the self back from the they’s idle chatter and common sense. Uncanniness is soundless, and the call coming out of this reticence calls Da-sein back to the stillness of itself. The call is clearly not objectively present, but that does not mean it’s not there to be heard.

The “eminent, authentic disclosedness” here attested, the “reticent projecting oneself upon one’s ownmost beig-guilty which is ready for Angst”, we shall call resoluteness. Disclosedness is primordial truth which does not refer to “judgement” but to a fundamental existential. In resoluteness we see Da-sein in this existential, the truth which is most primordial because it is authentic.

It’s important to recall the nature of Da-sein as being-in-the-world so that authentic being a self is not misunderstood as a seclusion from the world. Resoluteness as authentic disclosedness is authentically being-in-the-world, bringing the self into its being together with things at hand and taking care of them. Resolute Da-sein frees itself for its world, letting others with it “be” in their ownmost potentiality-of-being and disclosing the potentiality in concern which leaps ahead and frees. From this arises authentic beingwith-one-another.

Da-sein is always already in irresoluteness, it is summoned out of being lost in the they by resoluteness, which concerns Da-sein with its potentiality-of-being. “Resolution…first discovers what is factically possible in such a wy that it grasps it as it is possible as one’s ownmost potentiality-of-being in the they.”

The spatiality of the there is grounded in disclosedness, and situation is likewise grounded in resoluteness. “Situation is the there disclosed in resoluteness”. The call of conscience summons us to our potentiality-of-being, calls forth to the situation. “Resolute, Da-sein is already acting.”

Oy...

The Authentic Potentiality-for-Being-a-Whole of Da-sein, and Temporality as the Ontological Meaning of Care

0.61 Preliminary Sketch of the Methodical Step from Outlining the Authentic Being-a-Whole of Da-sein to the Phenomenal Exposition of Temporality

Authentic being-toward-death is anticipation, and authentic potentiality-of-being is resoluteness. The only way to bring the concepts together is to project resoluteness upon the most extreme possibility—to see if resoluteness points to anticipatory resoluteness as its ownmost authentic possibility.

To tackle this we must reconsider Da-sein, clarifying the phenomenon of care by figuring the role of “self” as neither substance nor subject in “the existng self whose being was conceived s care”. This leads to the exposition of temporality, which temporalizes itself in various possibilities of existence and which plays a role in anticipatory resoluteness.

We’ll do our usual bit to distinguish temporality from the vulgar “time”. Okay, let’s do it.

0.62 The Existentielly Authentic Potentiality-for-Being-a-Whole of Da-sein as Anticipatory Resoluteness

s anticipatory resoluteness a new phenomenon or is it an existentiell modalization of resoluteness through being-toward-death? Resoluteness is reticent self-projection upon one’s ownmost being-guilty, also demanding Angst of oneself. This being-guilty means a null being the ground of a nullity. This is not quanitifiable or intermittent, but constant. Only if resoluteness is transparent enough to understand it as such can “guilt” occur authentically in resoluteness. Resoluteness develops authentically as anticipation of death, or beingtoward-the-end-that-understands. Resoluteness and anticipation are not merely “connected” but intertwined: resoluteness has authentic being-toward-death as the possible existentiell modality of its own authenticity.

The being of Da-sein is primarily potentiality-of-being. because being-guilty belongs to the being of Dasein, “guilty is always only the actual factical potentiality-of-being.” Being-guilty must thus be conceived as a potentiality-for-being-guilty. Resoluteness understands itself in this potentiality-of-being, staying in a primordial possibility of Da-sein. It does so authentically only when it “is primordially what it tends to be.” This is being-toward-death, anticipation. “Resoluteness becomes a primordial being toward the ownmost potentiality-of-being of Da-sein only as anticipatory.” Anticipatory resoluteness is the only way to understand the potentiality-for-being-guilty authentically and wholly, i.e.: primordially.

Anticipation of death is disclosed as the nonrelational possibility. This bypasses “worldly” status. Wanting-to-have-a-conscience signifies readiness for the summons to one’s ownmost being-guilty that always already determined factical Da-sein before and after any factical indebtedness.

In its death, Da-sein must “take itself back”. Anticipating, i.e.: constantly being-certain of this, resoluteness is authentically and wholly certain. It must look to itself prior to relatedness in order to look to the possibility of nonexistence. Resoluteness looks to the whole potentiality-of-being of Da-sein. But even in this mode Da-sein holds itself open for the irresoluteness of the they. Resoluteness is forced to look to its potentiality-of-being with regard to the actual situation. The indefiniteness of death discloses itself primordially in Angst, an attunement that strives for resoluteness. We can no longer conceive of anticipation as hypothetical and unattached behavior. Anticipation is “the possibility of the authenticity of that resoluteness existentielly attested to in such resoluteness—a possibility concealed and thus also attested.”

So basically anticipatory resoluteness is a mode of being in which Da-sein is aware of its possibilities and how they all lead to death, specifically in a way that is not concerned with the details with the they’s everyday, commonsense interpretation.

0.63 The Hermeneutical Situation at Which We Have Arrived for Interpreting the Meaning of Being of Care, and the Methodical Character of the Existential Analytic in General

We have now defined Da-sein well enough in its ontological peculiarities, with enough primordiality, that we can properly interpret the meaning of the being of care. We have thus far shown how the commonsense way of taking care of things has taken over and closed off Da-sein’s potentiality-of-being.

The kind of being of Da-sein therefore needs any ontological interpretation to “be in charge of the being of this being in spite of this being’s own tendency to cover things over.” In circumspect taking care of things, one circumspectly heeds taking care of things as well. In existing and understanding, Da-sein somehow (appropriately or not) understands its own existence as well. In this way every ontological question Da-sein asks about being is made ready by the kind of being of Da-sein.

Now we confront our “circularity”. Does Da-sein’ interpretation of its existence inherit its guidelines from a “presupposed” idea of existence? This “circle” we are so well trained to avoid is in fact the basic structure of care! Da-sein has always already projected itself upon its possibilities, and in these projects it also has pre-ontologically projected “something like existence and being.”

But common sense only takes care of “factual” beings, beings in view of its circumspection. It ignores and fails to understand that “factual” experience comes after and along with understanding of being. Common sense brands anything outside its understanding of being. Common sense brands anything outside its understanding as “violent” because it misunderstands understanding. The being of Da-sein is circular, so rather than reducing it to a problem of an “I” and an object, we must leap into the “circle” primordially, letting ourselves proceed with a complete view of Da-sein’s circular being. Now to tackle care!

0.64 Care and Selfhood

To clarify care we must show it as Da-sein turns to itself in care. This begs a proper ontological examination of selfhood.

The everyday self-interpretation of Da-sein expresses “itself” in saying-I. This simple declaration refers to nothing but me, an absolute “subject”. This is a single persistent thing to which Kant gave the characteristics “simplicity”, “substantiality”, and “personality”. Time to find out where this comes from.

Kant pegs the phenomenal content of the “I” as “I think”, which is not something represented but rather “the formal structure of representing as such”, making possible all representations. This ties the I into an objectively present subject, leading back to the res cogitans. Kant’s empiricism holds him back from the phenomenological understanding of “I think” as “I think something”. Though he recognizes the dependency of the I upon its representations, he fails to discuss their connection (which is what’s actually there!).

In clarifying the “something”, we find it must be innerworldly, which implies a world. This shifts the I to an “I-am-in-a-world”. Everyday saying-I does not see itself in relation to its own kind of being, but rather tends to understand itself in terms of the “world” taken care of.

Selfhood can only be found existentially in authentic potentiality-of-being-a-self, in authentic Da-sein as care. Now we can understand, in terms of care, the constancy of the self as its having gained a stand. “Existentially, the constancy of the self means nothing other than anticipatory resoluteness.” In reticence, Da-sein “is” rather than repeatedly saying “I”.

Now, temporality.

0.65 Temporality as the Ontological Meaning of Care

“Meaning signifies…that in terms of which something can be conceived in its possibility as what it3is. Projecting discloses…what makes something possible.” To set forth the meaning of care is to pursue the project guiding Da-sein’s existential interpretation so tha its upon-which is made visible in what is projected.

“Meaning signifies the upon-which of the primary project of the understanding of being.” That is, meaning is at the root of the understanding of being. For a thing to “have meaning” is for it to have become accessible in its being. It “has” meaning because it becomes intelligible in the project of that being.

“The primary project of the understanding of being ‘gives’ meaning.”

Anticipatory resoluteness is the being toward one’s ownmost, eminent potentiality-of-being. This is possible as Da-sein is able to come toward itself in its ownmost possibility. Letting-come-toward-itself that “perdures the eminent possibility” is the primordial phenomenon of the future. This “future” isn’t what is “not yet actual”, but is how Da-sein comes toward itself in its ownmost potentiality-of-being. Since anticipation makes Da-sein authentically futural, anticipation is possible only in how Da-sein always already comes toward itself—Da-sein is futural in its being in general.

Da-sein is as I-am-having-been. “Authentically futural, Da-sein is authentically having-been.” Letting what is at hand be encountered in action can only happen in a “making that being present.” The present is released from the future that is in the process of having-been. We cann the unified phenomenon of this process temporality, and it is the meaning of authentic care, given this idea of making present.

So now we get to demonstrate how the vulgar, derivative concept of time arose from primordial and authentic temporality, which requires a concrete description of the latter phenomenon.

Temporality is where the primordial unity of the structure of care is. “Ahead-of-itself-already-in-a-world.” The ahead and already express temporality, but clearly not in the vulgar sense. “Ahead” is the future that first makes possible Da-sein’s ability to be concerned about its potentiality-of-being. “Already” is already having-been, meaning “I-am-as-having-been.” Da-sein, in attunement, finds that it is what it still is and already was.

But there’s another part to care: entangled being-together-with-. . .. Making present is the primary basis of falling prey to things at hand and objectively present. It is included in the future and in having-been, via primordial temporality.

“Future, having-been, and present show the phenomenal characteristics of ‘toward itself’, ‘back to’, ‘letting something be encountered’.” These reveal temporality as the “ekstatikon par excellence”, the primordial “outside of itself” in and for itself. The phenomena of future, having-been, and present are the ecstasies of temporality. Temporality temporalizes in these ecstasies’ utility, not through their cumulative sequence. Nevertheless, the future has priority in this unity. Authentic future, futurally having-been, arouses the present. Primordial temporality is thus finite, since it is related to anticipatory resoluteness. Primordial and athentic futuer is toward oneself, and says nothing of anything once no longer here. Inauthentic temporality temporalizes an infinite time out of finite time.

Now we have to relate this more concretely to Da-sein.

0.66 The Temporality of Da-sein and the Tasks Arising from It of a More Primordial Retrieve of the Existential Analysis

Time to reveal the temporal meaning of everydayness. “Temporality reveals itself as the historicity of Dasein.” In circumspect taking care of things, Da-sein expends itself, “uses itself up”. Da-sein uses its time, reckoning with it.

Circumspect and reckoning taking care of things discovers time and builds a measurement. Circumspect care lets what is at hand and objectively present be encountered in time, giving them a “within-time-ness”. The vulgar and traditional concept of time arises from this, and thus is grounded in and arises from primordial temporality. We have elaborated Da-sein’s temporality as everydayness, historicity, and within-time-ness. This should demonstrate the importance of clarifying beings unlike Da-sein for the ontological problematic. Everyday temporality will need much more clarification than this.

Temporality and Everydayness

0.67 The basic Content of the Existential Constitution of Da-sein, and the Preliminary Sketch of Its Temporal Interpretation

Primordial constitution of Da-sein is not a reduction to some “ultimate structural element”. Things “arise” in ontology by means of degeneration, not revealing something ontically self-evident by “common understanding” but instead calling all self-evidence into question.

Blah blah I’m pretty sure this is recap but we’ll see.

0.68 The Temporality of Disclosedness in General

The individual structural moments of care ar eunderstanding, attunement, entanglement, and discourse. To temporally interpret these moments is to demonstrate the concrete temporal constitution of care. Okay, so...

“Every understanding has its mood. Every attunement understands. Attuned understanding has the characteristic of entanglement. Entangled, attuned understanding articulates itself with regard to its intelligibility in discourse.” Looking at these we can reconstruct the “one temporality that holds within itself the possible structural unity of” these.

a. The Temporality of Understanding— The primordial and existential meaning of understanding: “to be projecting toward a potentiality-of-being for the sake of which Da-sein always exists.”

Understanding is the basis upon which Da-sein can develop the possibilities of sight, looking around, looking. Da-sein’s potentiality-of-being is disclosed to itself such that it knows its situation, what’s going on with itself in terms of existentiell possibility. Understanding oneself projectively happens because of the future: it is a coming-toward-oneself from the possibility as which Da-sein always exists. “Understandingly, Da-sein always is as it can be.”

Temporality does not constantly temporalize itself out of authentic future, so the temporalizing of the future is changeable.

Anticipation is still the term to use because the future arises not from the present but from the inauthentic future. Da-sein is always factically ahead-of-itself, but inconstantly anticipatory. In the mode of inauthentic future Da-sein does not anticipate but instead it awaits its ownmost, nonrelational potentiality-of-being “heedfully in terms of that which what is taken care of produces or denies.” Heedful expectation is founded in awaiting and temporalizes itself authentically as anticipation.

In resoluteness, the present isn’t just acquierd from the “dispersion in what is taken care of” and kept, but is “held in the future and having-been.” This authentic present is called the Moment. This phenomenon does not reduce to the now. “In the Moment” nothing happens, but there we can encounter for the first time what can be “in a time” as objectively present.

Contrasting the authentic Moment is the inauthentic present, called making present. Even authenticpresent makes present, but inauthentic present merely does so; it is not “in the Moment”. “[I]nauthentic understanding projects its potentiality-of-being in terms of what can be taken care of,… it temporalizes itself in terms of making present. The Moment, on the other hand, tmproalizes itself out of the authentic future.”

“In anticipation, Da-sein brings itself forth again to its ownmost potentiality-of-being. We call authentic having-been retrieve.” Inauthentic projecting is based on Da-sein having forgotten itself in its ownmost, thrown potentiality-of-being. This forgetting is not failure to remember, but a “ ‘positive’, ecstatic mode of having-been”. On the basis of this forgetting mere making present retains things, beings unlike Da-sein encountered in the world.

Just as awaiting makes possible expectation, forgetting makes possible remembering, and not the other way around.

“Although inauthentic heedful understanding is determined in the light of making present what is taken care of, the temporalizing of understanding comes about primarily in the future.”

b. The Temporality of Attunement— Understanding is always attuned. Attunement brings Da-sein before its thrownness, disclosing it in “how one is”. Being thrown, existentially, is to find oneself in such and such a way. Understanding is primarily grounded in the future, but attunement temporalizes itself primarily in having-been.

Ontically, moods come and go, running their courses “in time”, but we need to explore the existential and temporal constitution of mood, which is that it is a “bringing back to …”. We must show that moods can only “signify” existentielly on the basis of temporality. Now, the temporality of fear:

Fear is inauthentic attunement, fear of something innerworldly, a circumspect “letting something-come-toward-oneself”. Doesn’t this primarily have to do with the future? Fear is futural, and “related” to “something futural”, but is not possible on the basis of this alone. “[T]he awaiting that fears is afraid ‘for itself’,… fear of is a fearing about.” Fear confuses, causing one to forget and back away from a factical, resolute potentiality-of-being into “those possibilities of self-preservation and evasion that have already been circumspectly discovered beforehand.”

How about Angst? Angst brings Da-sein before its thrownness and the uncanniness of everyday, familiar being-in-the-world. The “threat” revealed in Angst is that everything at hand and objectively present has nothing to “say” to us. One has Angst in the face of Da-sein itself. In Angst one cannot project oneself upon a potentiality-of-being based on what is taken care of. Angst is anxious about Fa-sein’s thrownness into uncanniness, bringing one back to the pure That of this ownmost thrownness. This bringing back to thrownness is as something to be possible retrieved. This reveals the possibility of authentic potentiality-of-being that, as something futural in retrieving, comes back to thrown There. So the attunement of having-been that constitutes Angst has the specific ecstatic mode of bringing before the possibility of retrieval. Both modes of attunement, Angst and fear, are primarily grounded in having-been, but “Angst arises from the future of resoluteness, while fear arises from the lost present of which fear is fearfully apprehensive,…”. Even hope, seemingly futural, is grounded in having-been in that it brings relief from a burden. Ontologically, only attuned beings can be affected, such that in making present Da-sein can be brought back to itself as having-been. How stimulation and touching of senses in the merely alive are to be ontologically defined is another problem.

c. The Temporality of Falling Prey— The future primarily makes understanding possible, having-been makes mood possible, and the present makes falling prey possible. We interpreted falling prey before as idle chatter, curiosity, and ambiguity. We can most easily see falling prey’s temporality in curiosity, so we’ll focus on that.

Curiosity is a tendency by which Da-sein takes care of potentiality of seeing. This seeing, really a broader perceiving, lets what is at hand and objectively present be “bodily” encountered with regard to outward appearance. This is grounded in a present. This present gives the ecstatic horizon by which beings can be bodily present. Criosity does not make present to understand, but to see and have seen. This greed for the new attempts to withdraw from awaiting. Thus inauthentically futural, curiosity does not await possibility but instead “desires possibility as something real.”

This making present, awaiting definite, grasped possibility, makes possible ontologically curiosity’s notstaying. Awaiting is ecstatically modified into “arising”, pursuing making present.

When awaiting is thus modified ecstatically by pursuit, that modification is “the existential and temporal condition of the possibility of dispersion.”

Making present thus modified “makes present for the sake of the present.” Dispersed not-staying becomes the inability to stay at all. “The Moment brings existence to the situation and discloses the authentic ‘There’.” In constrast, this present does the opposite. It is an inability by which Da-sein “is everywhere and nowhere.”

ne forgets as the present “arises”. Curiosity has forgotten what went before and keeps to what is nearest by. This fact is the ontological condition for curiosity, not just a result from curiosity.

Da-sein, awaiting and forgetting, still is temporal in that it understands itself, albeit alienated from its ownmost potentiality-of-being grounded in authentic future and having-been. Curiosity, offering something “new”, does not let Da-sein come back to itself. Curiosity is not “brought about” by “the endless immensity of what has not yet been seen” but by the entanglement of the arising present.

The present arises from a fleeing from thrown being-toward-death, letting Da-sein come to authentic existence only by taking a detour through the present. The origin of entanglement is “the primordial, authentic temporality itself that makes possible thrown being-toward-death.”

Da-sein is swept along in thrownness. The present, existentially constituting being swept along, cannot move beyond itself into another ecstatic horizon unless brought back by resoluteness such that the actual situation and the “boundary situation” of being-toward-death are brought together as the held Moment.

d. The Temporality of Discourse— We cannot show the tmeproality of discourse until the fundamental connection between being and truth has been unfolded in the temporal problematic. Then we can find the ontological meaning of the “is”.

Understanding is grounded in the future (anticipation, awaiting), yet is always a present that “has-been”. Attunement temporalizes itself primarily in having-been (retrieve, forgottenness), yet also as a future that “makes present”. Falling prey is temporally rooted primarily in the present (making present, the Moment), yet “arises” from a future that has-been. The unity of the structure of care is thus grounded in the exstatic unity of the temporalizing of temporality. A future that makes present, in the process of having-been. Now we must show the ontological possibility of being-in-the-world.

0.69 The Temporality of Being-in-the-World and the Problem of the Transcendence of the World

“The ecstatic unity of temporality…is the confition of the possibility that there can be a being that exists as its ‘There’.” This is the clearing of Da-sein, clearedness arising from care. Before any “temporal” interpretation, care “opens” and makes Da-sein “bright” for itself. All the essential existential structures of Da-sein are regulated and unified by ecstatic temporality.

We have traced the origins of care to temporality, but it is now time to explicitly conceive taking care in terms of temporality.

We face the question: how is something like world possible? In what sense is world? What and how does it transcend, and how ar einnerworldly beings “connected” with the surrounding world? We’ll do this in three parts: tmeporality of circumspect taking care, the modification of this into theoretical knowledge of objectively present innerworldly things, and transcendence of the world.

a. The Temporality of Circumspect Taking Care— Heedful being together with the “world” is our association in and with the surrounding world, conceived earlier as being together with things that belong to everyday need. Taking care and things taken care of aren’t just objectively present together, but their connection is informative. What i taken care of can shed light on taking care. The kind of being of useful things at hand is relevance. That relevance is relational highlights that there is no such thing, ontologically, as a useful thing. Taking care that circumspectly discovers is letting something be in relevance. Letting things be relevant is taking care, which belongs to the essential constitution of care, which is grounded in temporality.

Thus, the possibility of taking care must be sought in temporalized temporality. “Understanding the intention and context of relevance has the temporal structure of awaiting.” The context is awaited and the relevance is retained, making possible the handy way in which a particular useful thing is made present. Being relevant constitutes itself in this way such that it makes absorption in taking care of a world possible. Useful things are not encountered through perception but in inconspicuousness.

The conspicuousness of, say, a broken took can not be discovered by “representation”, since it can be conspicuous only in and for the user. This damage “holds up” the making present that awaits and retains, shifting it more to itself as it works to inspect and remove the disturbance.

The disclosure of obstcles on the basis of the ecstatic temporality of taking care allows Da-sein to “understand itself in its abandonment to a ‘world’ of which it never becomes master.” Da-sein exists in a world because of a retention that awaits, which is why it in a way always already knows its way around even a strange “world”.

Now we’re going to demonstrate how theoretical behavior to the “world” comes out of circumspect taking care.

b. The Temporal Meaning of the Way in which Circumspect Taking Care Becomes Modified into the Theoretical Discovery of Things Objectively Present in the World— How does theoretical discovery “arise” from circumspect taking care? This is not about denying “the ontic history and development of science, or of its factical occasions or of its nearest goals”, but searching for the ontological genesis of theoretical behavior. What makes this mode of behavior possible? We seek not a “logical” conception of science (“a ‘context of causal relations of true, that is, valid propositions’ ”), but an existential concept of science as a mode of existence, a mode of being-in-the-world “which discovers or discloses beings or being.”

Science’s goal of pure observation seems at first to imply that taking care astains from use. But it is not simply a privation, the absence of praxis. This is just-looking-around, not a “theoretical” attitude. Scientific behavior is in no way a “purely intellectual activity”, but is regulated by “seeing”. We thus have to characterize the circumspection of the “practical” mode. There is a potentiality-of-being of Da-sein for the sake of which taking care exists as care. This potentiality-of-being illuminates the overview, the primary understanding of the totality of relevance. Circumspection brings things nearer to Da-sein by using in the mode of interpretation. We call this bringing near deliberation. This works as an “if-then”. Bringing near in circumspect deliberation does not simply “confirm” a being’s objective presence; it has the existential meaning of making present.

In awaiting a possibility, Da-sein takes care of useful things. Circumspect making present is grounded in a retention of the context of these useful things. Da-sein, awaiting a possibility (what-for), retains a thing at hand (for-this). Conversely can the making present belonging to awaiting retention bring it explicitly nearer in reference to the what-for. The deliberation that brings near must adapt to the kind of being of what is brought near.

The grounding of the present in the future and the having-been makes possible bringing near in a making present such that the present adapts itself to what is encountered, interpreting itself in the schema of the as-structure. The “as” is grounded in the unity of temporality.

The theoretical mode removes an object from deliberation. A hammer that is too heavy is now just “heavy”, describing a “property” of its “mass”. The hammer is no longer a tool and is now only a corporeal thing. It no longer provides us with a relation that can produce “too heavy”.

Switching to looking at the hammer as objectively presen, the understanding of being guiding heedful association has been transformed. This overlooks the place of the useful thing, turning it into a “location”, “world point”—a position in space and time. “Objectifying being together with things objectively present has the character of an eminent making present.” This discovering awaits the discoveredness of only what is objectively present.

The scientific project of nature requires the thematization of what is objectively present. This requires that Da-sein transcend the beings thematized. For thematization to modify the understanding of being presupposed an understanding of something like being. Handiness and objective presence aren’t differentiated, but there is still some unthematic relation to a context of useful things. “A world must be disclosed to it.” Temporality is the grounding of the being of Da-sein, and thus makes possible being-in-the-world and the transcendence of Da-sein. This supports the being together with innerworldly beings that takes care theoretically or practically.

c. The Temporal Problem of the Transcendence of the World— Taking care is grounded in in-order-to, what-for, for-that, and for-the-sake-of-which, which together constitute world. But how must world be in order for Da-sein to exist as being-in-the-world? The existential possibility of the world is possible because temporality has a horizon. Each ecstasy has a “whereto” of raptness, a horizonal schema.

• Da-sein comes back to itself futurally in the schema of for-the-sake-of-itself.

• The horizonal structure of having-been is characterized by the schema of that in the face of which Da-sein has been thrown and to which Da-sein has been delivered over.

• The horizonal schema of the present is determined by the in-order-to.

On the basis of the horizonal unity of these schemata of the ecstasies we can see the primordial connection between the in-order-to and the for-the-sake-of-which. Thus the being that is always its There has something like a disclosed world on the basis of the horizonal constitution of the ecstatic unity of temporality.

Neither objectively present nor at hand, the world temporalizes itself in temporality. “It ‘is’ ‘there’ together with the outside-itself of the ecstasies”, and is not “there” unless Da-sein exists and temporalizes itself.

In being together with things heedfully and factically, the world is already presupposed. Innerworldly beings can only be discovered in the There of Da-sein’s existence, and Da-sein’s control is limited to what, which direction, to what extent, and how it discovers and discloses.

The world is “further outside” than any object. “The relations of significance that determine the structure of the world are thus not a network of forms that is imposed upon some material by a worldless subject. Rather, factical Da-sein, ecstatically understanding itself and its world in the unity of the There, comes back from these horizons to the beings encountered in them.”

The “problem of transcendence” is not an escape from a subject to objects constituting a world, but the question of how beings encountered in the world become objectified as encountered beings.

Developing the structure of world relies upon “an ontology of possible innerworldly beings…oriented toward a clarified idea of being in general”. First we must set forth the temporality of Da-sein.

0.70 The Temporality of the Spatiality Characteristic of Da-sein

“…Da-sein must be addressed coordinately as ‘temporal’ ‘and also’ as spatial.” Temporality “embraces” spatiality as an existential foundation.

Spatiality is possible for Da-sein only as care—factically entangled existing. Da-sein does not fill out space as though objectively present, but rather Da-sein takes space in. Da-sein makes room for space, then comes back to a “place” it has taken over. It has made room for a leeway.

This making room is constituted by directionality and de-distancing. How is this grounded in temporality? Making room is what allows Da-sein to directionally discover something like a region. This region has already been discovered in the discovery and handling of useful things. “Being-in-the-world that takes care of things is directed, directing itself.. . . The self-directive discovering of a region is grounded in an ecstatically retentive awaiting of the possible hither and whither.” Making room is a directed awaiting of region, and is equiprimordially a de-distancing of things at hand and objectively present. This is “… grounded in a making-present that belongs to the unity of temporality in which dierctionality is possible, too.

The here of Da-sein’s factical situation signifies “the leeway of the range of the totality of useful things taken care of nearby”. This leeway is the opening made by directionality and de-distancing.

In falling prey, Da-sein brings near in a way that is founded in “making present”. This bringing near forets the over there, creating the illusion that a thing is “initially” objectively present.

The priority of “spatial representations” in language comes from the kind of being of Da-sein, not from the space itself. Temporality, on whose basis Da-sein is capable of breaking into space at all, meets up with these spatial relations asit gets lost in making present and uses them as guidelines for articulating understanding.

0.71 The Temporal Meaning of the Everydayness of Da-sein

Every structure of the constitution of being of Da-sein that we analyzed so that we could arrive at temporality must now be taken back and understood in tems of temporality. We started our analysis with Da-sein’s inconspicuous, average modes of existing, in which Da-sein holds itself initially and for the most part—everydayness. “Initially” signifies the priority of being-with-one-another in publicness, even if this is existentially “overcome”. “For the most part” means that Da-sein shows itself and exists in everydayness “as a rule”, but clearly not always. Everydayness is a being comfortable in habit, a stretching along “temporally” in the succession of days. This habit, this “like yesterday, so today and tomorrow”, and the “for the most part” require that we understand the “temporal” stretching along of Da-sein.

Basically we now see that everydayness means temporality, so back we come to having to understand being.

Temporality and Historicity

0.72 Existential and Ontological Exposition of the Problem of History

More primordial than the project of Da-sein’s authentic existence as a basis for understanding Da-sein is to not limit the investigation to futurality. Death is merely Da-sein’s “end”, and one such end at that. The other “end” of Da-sein is its “beginning”. Thus far we have examined Da-sein leaving “behind” everything that has been. The whole which ought to be the subject of our analysis is the “connection of life” in which Da-sein stretches along between birth and death.

Da-sein doesn’t fill up a stretch “of life” by accumulating discrete and objectively present moments of reality; it stretches itself along—its own being is alread this stretching along. The “between” is already in the being of Da-sein.

Factical Da-sein is not something that was born in a past no longer real. Da-sein is born, is dying. To clarify the “connectedness of life” will require an approach with Da-sein’s temporal constitution as its horizon.

The movement of the stretched out stretching itself along is Da-sein’s occurrence. By understanding the structure of occurrence we can understand historicity ontologically. Understood objectively, history becomes the scientific object of historiography, and attitude athat is derivative of the basic phenomenon of history, which concerns the self-constancy of Da-sein that is a mode of its being “grounded in a specific temporalizing of temporality”. We must investigate the rootedness of historicity in temporality.

Rooted in care, historicity is either authentic or inauthentic. The nearest horizon of everydayness is Da-sein’s inauthentic historicity.

Okay, so p. 345 is mostly nonsense, I think, but he claims that historicity and within-time-ness are equiprimordial, thus justifying vulgar understanding in his usual wa and preparing for an analysis thereof.

0.73 The Vulgar Understanding of History and the Occurrence of Da-sein

We must disambiguate the vulgar interpretation of “history” and “historical”. There are four meanigs to examine:

1. Something that belongs to former events but is still objectively present “now”. Example: remains of a Greek temple.

2. A “connection” of events “past”, “present”, and “future”. Derivation of past, “development”. “Past” has no priority here.

3. “Human history.” Distinct from nature. “Spirit” and “culture” of humankind.

4. That which has been handed down. Inherited knowledge.

These all posit human being as the “subject” of events. Why is the past so important here? We must explore the temporal meaning of the “past”. What makes an “antiquity” “past” when it is still objectively present in the “present”? What is “past” about them? The world in which they were encountered as useful things at hand in heedful circumspection by Da-sein, a world which is no longer. What does itmean that the world no-longer-is? World is only in the mode of existing Da-sein. That is, factically as being-in-the-world.” Historical character of antiquities is grounded in the “past” of Da-sein to whose world that past belongs. This means only “past” Da-sein can be historical, by which we mean Da-sein that has-been-there. Why does having-been predominately determine what is historical when it temporalizes itself equiprimordially with present and future?

Da-sein is primarily historical. Beings unlike Da-sein encountered in the world, including nature, are secondarily historical. These are world-historical, which is the source of the vulgar concept of “world history”. What about historicity?

0.74 The Essential Constitution of Historicity

Da-sein always has its “history” because its being is constituted by historicity. This all rests on temporality, which we first explores through the authentic existence that is anticipatory resoluteness—confronting death as one’s potentiality-of-being in order to take over the being that Da-sein is in its thrownness. Da-sein comes back to itself in resoluteness. Resoluteness takes over Da-sein’s heritage as thrown and discloses to Da-sein its factical possibilities of authentic existing in terms of that heritage. Being free for death—clearing its schedule—gives Da-sein its goal and establishes in a possibility both inherited and chosen.

Da-sein is not “subject” to fate—it is fate. Da-sein, resolutely handing itself down, is disclosed being-in-the-world for the “coming” of “fortunate” circumstances and the “cruelty of chance”.

Fateful Da-sein exists as being-in-the-world in being-with others. Its occurrence is therefore an occurrencewith, which is determined by destiny. This is the occurrence of a community. Destiny is from being-with-one-another in the same world and in resoluteness for definite possibilities. The complete, authentic occurrence of Da-sein is constituted by Da-sein’s destiny in and with its “generation”.

Fatefulness is being historical in the ground of one’s existence, and this can only happen in a being constituted as care.

So... the others with whom Da-sein is in the world constitute a ground, a common occurrence, a destiny. This is the context for a being constituted as care to have a “history”, historicity, and for being fateful, to hand itself down to itself.

A being must be equiprimordially futural and having-been in order to hand down its inherited possibility t itself, “take over its own thrownness and be in the Moment for ‘its time’.”

When resolute Da-sein hands itself down, and Da-sein goes back to the possibilities of Da-sein that has been there, this explicit handing down is Retrieve.

History as a mode of being of Da-sein has its roots in the future.

Retrieve hands down fateful destiny along with heritage. It is the handing down of project, goal, of one’s end.

All of this historicity is in corformity with the occurrence lying in anticipatory resoluteness, so we’ll call it Da-sein’s authentic historicity. This whole model still comes off as discrete, though, with Da-sein having successive resolutions and experiences. Where is the “connection”? Time to look at the origin of the question of the constitution of the connection of Da-sein. Perhaps our inquiry thus far has determined the question and blocked out our access to the “connection” of authentic historicity.

0.75 The Historicity of Da-sein and World History

Da-sein initially and for the most part understands itself in terms of what it encounters in the world and circumspectly takes care of. This understanding is a projection upon possibilities, so understanding things encountered and taken care of is to concern oneself not just with useful things but with what “is going on” with them.

“The occurrence of history is the occurrence of being-in-the-world.” This includes the useful things and objectively present things encountered by being-in-the-world, for they have their fates and histories. These beings are world-historical. Factical Da-sein is entangles, and so in understanding itself in terms of the things it takes care of it understands its history as world history. Everyday Da-sein is dispersed in the multiplicity of its affairs, of what it takes care of. To come to itself, it “must first pull itself together from the dispersion and disconnectedness of what has just ‘happened’ ”. Now we can glean from the horizon of inauthentic historicity just what the question of “connectedness” is.

We inauthentically see that Da-sein pulls itself together into its connection. How did it lose itself in the first place that it must be pulled together? Resolute, authentic Da-sein stands against the inconstancy of dispersion as a steadiness that has been stretched along. Resoluteness is existence’s loyalty to its own self. It is not an “act” that is performed sometimes but not others. Resoluteness provides the existentiell constancy which has anticipated every possible Moment. Resoluteness, as fate, frees Da-sein to give up a resolution. So constancy is not the joining together of Moments, but Moments arise from the already stretched along temporality of that retrieve which is futurally having-been.

It makes sense now to look for an ontological genesis of historiography as a science in terms of Da-sein’s historicity.

0.76 The Existential Origin of Historiography from the Historicity of Da-sein

Historiographical disclosure of history is rooted in Da-sein’s historicity, having the same ontological structure. Historiographical thematization of history can only occur following the disclosure of the “past” in general. This is only possible through Da-sein, so “what historiographical thematization presents as the possible object of its investigation must have the kind of being of Da-sein that has been there”.

World history is together with Da-sein as being-in-the-world. Once Da-sein is no longer there, the world is also something that has-been-there. World-historical objectively present things can become historiographical material by being understood by Da-sein that has-been-there. This doesn’t return to the “past”; it presupposed historical being toward Da-sein that has-been-there, the historicity of the historian’s existence. This existence is what existentially grounds historiography as a science. From here we can determine the object of historiography, which must be linked to authentic history and its disclosure of what-has-been-there. This is retrieve. One cannot have authentic historiographical understanding, however, if the possibility to be retrieved is distorted into notions of “individual” events and universal “laws”.

Arising from the historicity of Da-sein, historiographical disclosure does not work backwards from a “present” to a “past”, but temporalizes itself out of the future, “selecting” its objects along with Da-sein’s existentiell choice, or rather grounded in it.

Remember that historiography’s central theme is not simply existence that has-been-there but its possibility. This existence always factically exists world-historically, so historiography is oriented toward “facts”.

Suddenly “we” are biting off Nietzsche! Heidegger’s hot it down that historiography isn’t necessary for historicity, but Nietzsche asks whether it’s beneficial or detrimental to have historiography and so describes three kinds, which we shall Marty up:

Monumental : Historiography arising from historicity that resolutely comes back to itsef, opening itself to the possibility of retrieve for the “monumental” possibilities of human existence.

Antiquarian : When historicity steps beyond retrieve and grasps the possibility to reverently preserve what has-been-there which made manifest the possibility to be retrieved.

Critical : When authentic, monumental-antiquarian historiography ceases to make present the entangled publicness o today and exists as a critique of the “present”.

So this “threefold character of historiography” is prefigured in Da-sein’s historicity. Authentic historicity makes possible their unity, but this all comes back to “temporality as the existential meaning of being of care.”

Time to see how this all works. OK, let’s do it...

0.77 The Connection of the Foregoing Exposition of the Problem of Historicity with the Investigations of Dilthey and the Ideas of Count Yorck

Dilthey and Count Yorck! We are talkin’ about history, here. Dude. Yorck is pretty sweet, all on about the “generic difference between the ontic and the historical.” So what we gotta do now is find a more primordial unity of the “ontic” and the “historiographical”.

So, what we gotta do is see to these points:

1. The question of historicity is an ontological question about the constitution of being of historical beings.

2. The question of the ontic is the ontological question of the constitution of being of beings unlike Da-sein, of what is objectively present in the broadest sense.

3. The ontic is only one area of beings.

The “ontic” and the “historigraphical” are encompassed in the idea of being. Wait... does this mean we need to clarify fundamental ontology first?

NO

FUCKING

WAY!

Dude.

Temporality and Within-Timeness as the Origin of the Vulgar Concept of Time

0.78 The Incompleteness of the Foregoing Temporal Analysis of Da-sein

Prior to any thematic inuvestigation, such as the sciences of history and nature, is Da-sein “reckoning with time”. This precedes instruments like clocks, makes them possible. Before proceeding, we must clarify how Da-sein reckons with time—what does it mean that beings are “in time”?

Da-sein, as temporality, temporalizes itself in a mode of behavior related to time in such a way that it “takes accound of it.” The earlier interpretation is clearly primordial time, and from it we can find how Da-sein levels it down to come to the vulgar conception of time.

Before asking whether time has “being” and how it can be so designated, we must investigate the role of temporality in Da-sein’s understanding of being. Good news if you like Hegel!

0.79 The Temporality of Da-sein and Taking Care of Time

Circumspect taking care is “calculating, plamning, preparing ahead, and preventing”. It says “then” … that will happen, “before” … that will get settled, “now” … that will be made up for, that “on that former occasion” failed or eluded us.

Taking care expresses itself in…

“then” —awaiting

“on that former occasion” —retaining

“now” —making present

The “then” is “spoken in a making present that awaitingly retains or forgets”; the “now not yet” is contained within.

Likewise, the “on that former occasion” contains the “now no longer”, wherein retaining expresses itself as a making present that awaits. These are understoof with regard to a “now”, and making present carries a lot of weight.

Horizons of retaining:

“on that former occasion” —“earlier”

“then” —“later on”

“now” —“today”

But what are these as such?

“then” —“then, when …”

“on that former occasion” —“on that former occasion when …”

“now” —“now that …”

This is called datability. Why does Da-sein express these concepts? Because in so doing Da-sein also expresses itself, in that it is expressing circumspect and understanding being-together-with things at hand. This is necessarily grounded in a making-present. This interpretedness that takes care is how temporality is initially and for the most part made known.

“Then” also contains the “and now not yet”—the awaiting that makes present understands the “until then.” This is an in-between, expressed datably in the “meanwhile”.

This is all a “during” that makes present and awaits. The during is the time revealed in the self - interpretation of temporality, understood as a “span”. The making present that awaits and retains is thus disclosed to itself as being ecstatically stretched along in historical temporality.

Example: “now” in the intermission, at dinner; “then” at breakfast. The taking care that awaits, retains, and makes present “allows itself” time, giving itself time in taking care without determining it prior to any specific reckoning.

Time here dates itself in terms of what one does “all day long”. The mor eDa-sein is absorbed in awaiting what is taken care of, the less it awaits itself and Da-sein begins to forget itself. In so doing, the time Da-sein “allows” itself is covered over.

“We often cannot bring a ‘day’ together again when we come back to the time that we have ‘used’.” Because of this covering over, the time Da-sein allows itself has “gaps” in it. This doesn’t mean it goes to

pieces, but rather means that this time with gaps is a mode of temporality that is disclosed and ecstatically stretched along.

We can only appropriately phenomenally explicate this mode in which “allowed” time “elapses” if we “avoid the theoretical ‘representation’ of a continuous stream of nows” and instead consider the modes in which Da-sein allows itself time in terms of “how it ‘has’ its time in a manner corresponding to its actual existence.”

Irresolute, inauthentic existence temporalizes as making present that does not await but forgets, and in losing itself in what is taken car eof irresolute Da-sein also loses its time. Hence, “I have no time.”

Resolute Da-sein temporally exists as the Moment, fatefully whole and stretched along in the authentic, historical constancy of the self. The There is disclosed only as situation, and resolute Da-sein “constantly” has time for what the situation requires.

Several people saying “now” together are dating this “now” differently. Thus Da-sein is making the time it interprets and expresses public in that it is being-with as being-in-the-world.

Everyday taking care, understanding itself in terms of the “world” taken care of, knows its “time” not as its own, instead heedfully exploiting the time “there is”, with which the they reckons.

0.80 Time Taken Care of and Within-Timeness

Is public time “merely subjective”? “Objectively real”? Neither? First we must precisely determine the phenomenal character of public time.

In taking care, one orients oneself toward time, making it public, available for everyone. This “public time” is the time “in which” innerworldly things at hand and objectively present are encountered, beings unlike Da-sein. We’ll now call such beings within-time, which will help us understand “public time”.

Circumspect being-in-the-world requires the possibility of sight, a need which led to the primacy of day and night in our reckoning with time. “The occurrence of Da-sein is a daily one by reason of interpreting time by dating it—a way that is prefigured in its thrownness into the There. Temporality is simultaneously the condition for the discoverability of the clock and of the possibility of the clock’s factical necessity.

In taking care, time is always already understood as time for . . .. The “now that so and so” is as such either appropriate or inappropriate. Interpreted time has the character of “time for …” or “not the time for …”. Time made public, as time for …, has the nature of world, so we’ll call it world time. Not because it’s innerworldly (it can’t be), but because it belongs to the world.

Time taken care of “is datable, spanned, and public and, as having this structure, it belongs to the world itself.”

“Time” is neither objectively present in the “subject” nor in the “object”, neither “inside” nor “outside”. Temporality temporalizes world time that constitutes a within-timeness of things at hand and objectively

present. These beings are unlike Da-sein, and so they can not be called temporal since they are atemporal “whether they occur, arise, and pass away realistically, or subsist ‘ideally’.”

Everyday taking care that gives itself time finds “time” in innerworldly beings. Thus, our look at the genesis of vulgar time starts from within-timeness.

0.81 Within-Timeness and the Genesis of the Vulgar Concept of Time

What’s up with clocks? By following the clock hands in a way that makes present, one counts them. In making present the moving hands, one says “now here, now here, and so on”, counting nows. This world-time we’ll call now-time.

Lost in useful things taken care of, Da-sein finds the time belonging to these things. The more this is done, the more is said: now, then, on that former occasion, showing time vulgarly as a succession of objectively present nows “that pass away and arrive at the same time.”

These nows lack both datability and significance, levelled down so that they run along after one another to constitute a succession. More and mor ethe nows are “seen” as themselves objectively present, as though “beings are encountered and also the now.”

This is most penetratingly revealed in the vulgar interpretation’s main thesis—the “infinitude” of time. This is possible only on the basis of “an orientation toward an unattached in-itself of a course of nows objectively present.”

Heedful flight from death looks away from the end of being-in-the-world. Death is always my own, and so the they is unable to die, it “always has more time”. Public time, levelled down, belongs to everyone, and thus no one. Ecstatic, horizonal temporality temporalizes itself primarily from the future, but the vulgar understanding starts with the now, specifically the pure now we cal the “present”. There is some provenance and factical necessity of vulgar time, and something about Hegel and spirit, so here we go.

0.82 The Contrast of te Existential and Ontological Connection of Temporality, Da-sein, and World Time with Hegel’s Interpretation of the Relation between Time and Spirit

a: Hegel’s Concept of Time— For Hegel, time is the “intuited” becoming, an unthought transition between nows wherein being “goes over” to nonbeing and vice versa. This is clearly vulgar, and we’ll think of it as the negation of negation.

b: Hegel’s Interpretation of the Connection between Time and Spirit— For Hegel, spirit conceives itself, it is grasping the non-I. The essence of spirit is the negation of a negation.

Hegel sought “concretion” of the spirit, and we seek the “concretion” of factically thrown existence. Hegel sought to describe how “spirit” falls intdo time through the above connection and primordial kinship of the two, and we seek to describe how, falling prey, factical existence falls out of primordial, authentic temporality.

0.83 The Existential and Temporal Analytic of Da-sein and the Question of Fundamental Ontology as to the Meaning of Being in General

The being of Da-sein, and the distinction between this and beings unlike Da-sein, is merely the point of departure for the ontological problematic. We must understand the nature of the “reification” that constantly threatens investigation of the being of Da-sein.

We require a secure horizon for questions and answers. “The strife in relation to the interpretation of being cannot be settled because it has not yet even been started.”

Da-sein, thanks to our clarification of its being, can now be related to beings, those encountered in the world and that of itself in existing.

How is this understanding possible?

“The existential and ontological constitution of the totality of Da-sein is grounded in temporality. Accordingly, a primordial mode of temporalizing of ecstatic temporality itself must make the ecstatic project of being in general possible. How is this mode of temporalizing of temporality to be interpreted? Is there a way leading from primordial time to the meaning of being? Does time itself reveal itself as the horizon of being?

OK, let’s do it...